### v1

# Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets

Abstract.
Type inference is a crucial task for reusing online code snippets, often found on platforms like StackOverflow, which frequently lack essential type information such as fully qualified names (FQNs) and required libraries. Recent studies have leveraged Large Language Models (LLMs) to perform type inference for such code snippets, demonstrating promising performance. However, these results are potentially affected by a data leakage issue, as the benchmark suite (StatType-SO) used for evaluation has been publicly available on GitHub since 2017, with the full suite available since 2023. Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data.
To comprehensively assess the type inference capabilities of LLMs on Java code snippets and identify potential limitations of LLMs, we conducted a three-pronged evaluation. First, utilizing Thalia, a program synthesis technique, we created ThaliaType—a new, previously unreleased dataset designed for type inference evaluation. ThaliaType comprises 300 unseen code snippets, generated in a manner consistent with the setting of StatType-SO. Our results revealed that LLMs exhibit a significant drop in performance when generalizing to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall. To further investigate the limitations of LLMs in understanding the semantic meaning of the code, we developed semantic-preserving code transformations. Our analysis showed that LLMs performed significantly worse on code snippets that are syntactically different but semantically equivalent. Finally, we employed delta debugging to distill the minimal syntax elements in snippets sufficient for inference by LLMs. We observe that while the type inference task primarily involves inferring FQNs for types used in code snippets, LLMs could still correctly infer FQNs even when the types were absent from the snippets. This suggests that LLMs rely more on re-using knowledge from training data instead of thoroughly analyzing the snippets.
Our experiments suggest that the strong performance of LLMs in prior evaluations was likely influenced by data leakage in the benchmarks, rather than a genuine understanding of the semantics of code snippets. Our findings highlight the crucial need for carefully designed and rigorously evaluated benchmarks to accurately assess the true capabilities of LLMs for type inference tasks. To ensure reliable and rigorous evaluations of LLMs’ type inference capabilities, future research should prioritize the use of unseen benchmarks that are explicitly excluded from the LLMs’ training data.
1. Introduction
Java code snippets found online, such as those from StackOverflow, are often not compilable or directly reusable because they lack crucial type information, including fully qualified names (FQNs, i.e., import statements) and required libraries. Developers must manually infer and recover this missing information in order to reuse these code snippets (Terragni et al., 2016; Terragni and Salza, 2022). This process requires understanding the libraries and identifying the correct FQNs based on the names and functions used in the code snippets. For instance, if a code snippet references a type with the simple name Logger, the developer must determine whether the appropriate import is java.util.logging.Logger from Java’s utilities library or org.slf4j.Logger from another logging library, depending on the methods invoked on the objects of Logger in the code snippet.
Type inference for code snippets can assist developers in recovering missing type information needed to reuse these code snippets (Dong et al., 2022; Subramanian et al., 2014; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023b; Kabir et al., 2024). Existing approaches to type inference for Java code snippets generally fall into two categories: constraint-based and machine learning (ML)-based techniques. Constraint-based methods build a knowledge base of libraries and the types contained within (Dong et al., 2022; Subramanian et al., 2014). The state-of-the-art technique SnR (Dong et al., 2022) leverages this knowledge base and uses constraints extracted from the code snippet to identify the precise types being used. In contrast, ML-based approaches (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024) learn from prior examples of type usage in existing code to perform type inference.
ML-based techniques often produce incorrect inferences, due to their limited understanding of code structure and the rules governing Java’s type system. Recent advancements in large language models (LLMs), such as OpenAI’s proprietary GPT family and Meta’s open-weight Llama models, offer new possibilities for overcoming these limitations. By leveraging training on diverse and extensive datasets, LLMs have shown potential in performing various software engineering tasks (Hou et al., 2024). Prior studies (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) suggest that LLMs achieved performance comparable to that of the state-of-the-art, SnR.
However, the evaluation of the prior techniques (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) is potentially affected by data leakage (Inan et al., 2021) where the training dataset includes the benchmark dataset, and data leakage allows LLMs to regurgitate the training data rather than conducting type inference. The benchmark suite, StatType-SO, used for evaluating type inference techniques has been publicly available on GitHub since 2017 (Sta, 2017), with the full suite available since 2023 (Sta, 2023). As shown in Table 1, recent state-of-the-art LLMs have knowledge cutoff dates in late 2023, creating the potential for leakage. Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data. Since the exact training data for LLMs are kept confidential, detecting such leakage is difficult (Carlini et al., 2021; Nasr et al., 2023). While recent work has called attention to this data leakage problem for software engineering research (Ozkaya, 2023; Sallou et al., 2024), no study has thoroughly, empirically evaluated LLMs’ performance on type inference on code snippets.
In this paper, we thoroughly evaluate LLMs’ type inference performance on Java code snippets to better understand the strengths and limitations of LLMs on type inference tasks, and uncover potential limitations of LLMs. We organized our inquiry into the following three research questions.
RQ1: How well do LLMs perform type inference on unseen code snippets?
To address data leakage, we generated a new dataset named ThaliaType using Thalia, a program synthesis technique originally designed for testing static type systems. ThaliaType comprises 300 code snippets guaranteed to be unseen by the evaluated LLMs, thereby avoiding data leakage. Furthermore, we used the same set of libraries as used in StatType-SO to generate ThaliaType, to (1) enable a direct comparison of type inference performance between the two datasets, and (2) ensure the evaluated LLMs were consistently and sufficiently trained on the same set of types used in both datasets. By evaluating LLMs on ThaliaType and contrasting those results with those from StatType-SO, we can better understand the impact of data leakage on performance.
RQ2: To what extent do LLMs understand the semantics of code snippets?
To investigate whether LLMs truly understand the semantics of code snippets during type inference, we designed semantic-preserving code transformations to create syntactically different versions of the input code (section 5). These transformations were intended to hinder LLMs’ reliance on memorized training data by making direct recall more challenging. By observing how LLMs’ performance changes on these semantically equivalent but syntactically distinct versions, we can gain insights into their ability to grasp the underlying semantic meaning of the code.
RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
Lastly, we sought to distill the minimal syntax elements sufficient for LLMs’ inference through delta debugging (Zeller and Hildebrandt, 2002) (section 6). By minimizing the input code to the LLMs while maintaining the LLMs’ inference results, we assess whether LLMs rely on superficial patterns from training data rather than genuinely analyzing code semantics. If LLMs are truly performing type inference rather than leveraging memorized patterns, essential elements such as simple names should be retained in the minimized snippets.
Following the recommendation from prior software engineering research (Sallou et al., 2024), we evaluate LLM’s performance using two open-source models Llama3.1:8b and Llama3.1:70b, alongside state-of-the-art closed-source models GPT-4o and GPT-4o-mini (see Table 1(b)). As a baseline, we compare these models against SnR (Dong et al., 2022), the state-of-the-art constraint-based type inference technique. This evaluation aims to provide deeper insights into the potential limitations of using LLMs and to guide future research in evaluating LLMs’ performance on other software engineering tasks.
In RQ1, we found LLMs are highly likely to be affected by the data leakage issue and are prone to making incorrect inferences on unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall for Llama3.1:8b. On ThaliaType, even the best LLM GPT-4o only obtained a precision of 54.74% and a recall of 44.54% . On the other hand, SnR not affected by the data leakage issue, obtained a precision of 84.15% and a recall of 84.43% on ThaliaType. This stark discrepancy suggests LLMs are affected by data leakage, thus prompting further investigation into their type inference performance.
In RQ2, we found that LLMs performed significantly worse on transformed code snippets that were syntactically different but semantically equivalent. Applying all transformations caused a significant drop in F1 scores on the StatType-SO dataset across all LLMs. This suggests that LLMs rely heavily on the exact syntactic structure, indicating a lack of semantic understanding that limits their generalizability in type inference.
In RQ3, delta debugging revealed that LLMs can infer the same types using only 6.37% of the tokens in StatType-SO, with the number of unique simple names decreasing by 41.49% after minimization. Despite the type inference task primarily involves inferring FQNs for the types in the code snippet, LLMs would still correctly infer FQNs even when the types were absent. This suggests that StatType-SO may enable LLMs to perform well by exploiting superficial patterns in code snippets; this performance may be due to data leakage rather than a genuine understanding of code semantics.
Our experiments indicate that prior evaluations of LLMs were likely influenced by data leakage rather than a genuine comprehension of the semantics of the code snippets. To address these in the future, it is essential to evaluate LLMs on benchmarks that effectively measure their performance and are guaranteed to be excluded from the LLMs’ training data such as using Thalia to generate unseen code snippets.
Contribution. This paper makes the following contributions.
We proposed a new dataset ThaliaType along with an approach to creating such new datasets to rigorously evaluate type inference performance of LLMs while mitigating the potential biases introduced by data leakage. With the help of ThaliaType, we demonstrated that prior evaluations using StatType-SO are likely subjected to the data leakage problem and thus the performance results are inflated and misleading.
We investigated how well LLMs understand the semantics of code snippets by experimenting with whether semantic-preserving transformations on the code snippets affect the type inference performance of LLMs. The evaluation results on StatType-SO and ThaliaType show that the precision and recall of LLMs on type inference are dependent on the exact content in the code snippet despite preserving the semantic meaning of the code, indicating the LLMs lack a deep comprehension of the semantics and heavily rely on superficial syntactic patterns for type inference.
We further investigated the essence of LLMs performing type inference by isolating the syntactic elements sufficient for LLMs’ using delta debugging. The results show that some essential simple names are not required by LLMs for type inference especially on StatType-SO, indicating LLMs may rely more on exploiting superficial patterns from their training data than on the semantic information in code snippets.
This study provides a novel approach to rigorously assess LLMs’ comprehension of program semantics. To enable replication, reproduction, and further research on LLMs for code-related tasks, we have made our dataset and code publicly available at https://github.com/uw-pluverse/thalia-type.
2. Type Inference on Java Code Snippets
Type inference on code snippets aims to automatically determine the precise types within the code, specifically their FQNs (Subramanian et al., 2014; Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024). This process is essential for effectively reusing code snippets from sources like StackOverflow, where type information is often incomplete or missing entirely.
A key challenge in type inference lies in resolving ambiguities with simple names. For instance, as shown in Figure 2(a), the simple name Logger could refer to types defined in various libraries, such as the JDK (JDK, 2014), SLF4J (SLF, 2021), Log4j (Log, 2024a), or Logback (Log, 2024b). Disambiguating these references necessitates a deep semantic analysis of the snippet, including method signatures (e.g., getLogger), and their argument types. The complexity of this task is significantly increased by the vast search space of possible type combinations across numerous libraries, making the inference process computationally demanding.
This section provides an overview of state-of-the-art machine learning-based approaches to type inference, along with our prompting strategy (section 2.1). Additionally, we briefly introduce SnR, a constraint-based type inference technique, as a baseline for comparison against LLMs (section 2.2).
2.1. Machine Learning-Based Type Inference
Machine learning-based methods use statistical models and learned representations of code to conduct type inference and resolve ambiguous simple names in code snippets (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024). Of these, LLMs have recently achieved state-of-the-art performance on type inference on the StatType-SO (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) dataset. Existing approaches typically employ a simple but effective prompt to guide LLMs in generating FQNs for types in the code snippet (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024). Notably, Kabir et al. (2024) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt. This iterative process achieved a precision of 99.5% on the StatType-SO dataset. In contrast, our work focuses on the issue of data leakage in type inference rather than repairing code snippets. Data leakage is a fundamental problem that compromises the validity of results, regardless of whether a single-prompt or iterative approach is used. To ensure consistency and comparability with prior research, we adopted a single-prompt approach for type inference. Following established practices (Huang et al., 2023a; Kabir et al., 2024), we developed a straightforward but highly effective prompt for LLM-based type inference.
Prompt Template. Our prompt, shown in Figure 1, is designed to guide the LLM in inferring the type information in Java code snippets. It includes two key components: (1) a system message that sets the LLM’s role as a helpful assistant, and (2) a user instruction explicitly requesting the addition of import statements for Java code. The placeholder {input_code} is replaced with the actual code snippet during inference. Figure 1(a) presents the exact wording of the prompt, while Figure 1(b) demonstrates the corresponding output generated by the LLM for an example code snippet. Despite its simplicity, our prompt achieves high accuracy and aligns with best practices for LLM prompt engineering (LLM, 2024). Notably, our result using GPT-4o on the StatType-SO dataset represents a 7% improvement over the prior results (Kabir et al., 2024), demonstrating the robustness of our prompt while maintaining a similar level of complexity.
| System | You are a helpful programming assistant. |
| User | Add import statements to the following Java code. Do not use wildcard imports. Include only the necessary import statements. Do not import nonexistent types. Please note that you need to pay close attention and your response should be specific and accurate. Avoid repetition. Reply with the import statements only. |
| {input_code} |
2.2. Constraint-Based Type Inference with SnR
SnR (Dong et al., 2022) is the state-of-the-art technique that uses constraint-based type inference (Dong et al., 2022; Subramanian et al., 2014). It has proven highly effective for inferring types in code snippets from the StatType-SO dataset. To illustrate the information SnR leverages, consider the example code in Figure 2(a). SnR processes the snippet by extracting constraints, such as identifying a type with the simple name Logger, which has a getLogger method that accepts a java.lang.String argument. Using the knowledge base shown in Figure 2(c), SnR assigns java.util.logging.Logger to Logger, satisfying the constraints. This approach precisely eliminates incorrect candidates, such as org.slf4j.Logger, which lacks the required getLogger method.
SnR demonstrates how semantic relationships in the code snippet alone can solve type inference problems. By evaluating LLMs alongside SnR, we aim to analyze the extent to which LLMs leverage such semantic information for type inference. These findings will provide deeper insights into LLMs’ type inference capabilities and highlight opportunities for future research.
| Simple Name | Inferred FQN |
| String | java.lang.String |
| Logger | java.util.logger.Logger |
| Simple Name | FQN | Method Name | Return Type | Argument Types |
| Logger | java.util.logging.Logger | getLogger | java.util.logging.Logger | java.lang.String |
| Logger | java.util.logging.Logger | info | void | java.lang.String |
| Logger | org.slf4j.Logger | info | void | java.lang.String |
| Logger | org.slf4j.Logger | debug | void | java.lang.String |
| String | java.lang.String | length | int | |
3. Datasets
We utilize two datasets StatType-SO and ThaliaType for evaluating LLMs’ type inference performance on code snippets.
3.1. StatType-SO
The StatType-SO dataset is prevalent for testing type inference tools (Phan et al., 2018; Saifullah et al., 2020; Dong et al., 2022; Huang et al., 2023b). It is constructed from real-world, manually repaired Java code snippets from Stack Overflow. StatType-SO contains 267 code snippets from six popular Java libraries (Android, GWT, Hibernate, JDK, JodaTime, XStream). The snippets were extracted from 50 randomly selected Stack Overflow posts for each library. To prepare the dataset, all parsing errors in the code snippets were first corrected. Next, the missing import statements were manually inferred and added to the code snippet, serving as the ground truth for evaluation.
While StatType-SO is a widely adopted dataset for evaluating type inference techniques, it is important to acknowledge the potential for data leakage. Given its long-standing availability since 2017 at GitHub, StatType-SO may have been inadvertently incorporated into the training data of some LLMs, potentially leading to artificially inflated performance scores.
3.2. ThaliaType: A New Dataset
In this paper, we introduce ThaliaType, a new dataset specifically designed to rigorously evaluate the type inference capabilities of LLMs while mitigating the potential biases introduced by data leakage. ThaliaType is generated using Thalia, a tool originally developed for testing compilers’ type checkers. Recognizing its potential for LLM evaluation, we repurposed Thalia (Sotiropoulos et al., 2024) to generate a diverse set of syntactically valid, well-typed, and type-intensive Java programs. Given a set of types, along with the fields and methods defined for each type, Thalia generates programs that utilize a subset of these components. Each program consists of a single class with a main method containing variable declarations, field accesses, and method calls.
Thalia’s ability to generate syntactically correct and well-typed programs, combined with their type-intensive nature, aligns perfectly with the requirements for rigorous type inference evaluation. Thalia ensures that the programs are both valid and well-typed, eliminating the need for manual code cleanup or ground truth labeling. Furthermore, the type-intensive nature of these programs presents a meaningful challenge for type inference, while the inclusion of field accesses and method calls provides useful context for LLMs during inference.
To create ThaliaType, we used Thalia to generate 300 code snippets, with the types, fields, and methods drawn from the six libraries used in StatType-SO. For each library, 50 code snippets are generated. By selecting the same libraries, we ensured that the types in the generated code snippets were familiar to the LLMs, eliminating performance differences based on varying familiarity with the libraries. Figure 3 compares the statistics of ThaliaType and StatType-SO. The code snippets in both datasets are similar in terms of lines of code and the number of assignments. While Thalia-generated programs are type-intensive, ThaliaType includes fewer method calls but utilizes more types (via import statements) in each snippet.
Our experiments show that SnR achieved comparable performance on both ThaliaType and StatType-SO (section 4.2), demonstrating that ThaliaType provides sufficient semantic information for the type inference task. To the best of our knowledge, ThaliaType is the first dataset designed for evaluating LLMs on code snippet type inference while addressing data leakage. Future research can leverage our replication package to generate additional code snippets using Thalia using their desired libraries to evaluate newer state-of-the-art LLMs.
4. RQ1: How well do LLMs Perform Type Inference on Unseen Code Snippets
The data leakage problem is a real threat when evaluating LLM type inference performance, which motivated us to perform a thorough evaluation to uncover their true type inference capabilities. To mitigate this issue, in this RQ, we compared LLMs’ type inference performance on the publicly available StatType-SO dataset and the unseen code snippets in ThaliaType, providing insight into the potential impact of data leakage.
4.1. Method
Code snippets from StatType-SO and ThaliaType are given to LLMs using the prompt pattern in Figure 1 and also our baseline SnR one at a time without import statements. To check the output, import statements in LLMs’ response are extracted and compared against the original import statements in the code snippets. To ensure reproducibility, all the inferences are performed with a fixed seed (set to one) and a temperature of zero. GPT-4o-mini and GPT-4o are accessed using the OpenAI API (OPE, 2024b), while Llama3.1:8b and Llama3.1:70b are accessed through a self-hosted Ollama API (OLL, 2024) running on an A100 GPU.
The precision, recall, and F1-scores are computed to evaluate the performance of type inference. Precision measures the proportion of correctly inferred FQNs among all inferred FQNs, while recall measures the proportion of expected FQNs that are correctly inferred.
4.2. Results
LLMs achieved lower performance on unseen code snippets in ThaliaType despite their strong results in StatType-SO (Figure 4). Specifically, as shown in Figure 4(b), on the unseen ThaliaType code snippets, GPT-4o achieved a precision and recall of 54.74% and 44.54% respectively, a 42.8% and 53.1% decrease in precision and recall compared to results on StatType-SO. SnR outperformed all LLMs in ThaliaType with a precision of 84.15% and a recall of 84.43% . Even on StatType-SO, SnR achieved second best precision and recall of 95.50% and 91.46% . The large disparity between the LLMs’ performance on StatType-SO and unseen code snippets in ThaliaType suggests that LLMs are highly likely affected by data leakage which potentially inflated LLMs’ type inference performance on StatType-SO.
Unlike LLMs, SnR relies solely on the names of types, method names, and relationships between types in code snippets. The fact that LLMs outperformed SnR on StatType-SO suggests that LLMs leverage additional information not used by SnR to perform type inference. Conversely, the poor performance of LLMs on ThaliaType compared to SnR implies that LLMs failed to fully exploit the simple names, method calls, and type relations within the code snippet. These performance discrepancies motivated further investigations into the specific factors contributing to LLMs’ type inference capabilities. In particular, we explore in RQ2 (section 5) if LLMs leverage the available semantic information in code snippets or simply rely on matching exact syntax of the code snippets from training data.
| Document Frequency | [0,1e2) | [1e2,1e3) | [1e3,1e4) | [1e4,1e5) | ¿=1e5 | ||||||
| StatType-SO | Total FQNs | 27 | 69 | 312 | 429 | 463 | |||||
| TP | Recall | TP | Recall | TP | Recall | TP | Recall | TP | Recall | ||
| SnR | 27 | 100.00% | 60 | 86.96% | 297 | 95.19% | 372 | 86.71% | 433 | 93.52% | |
| Llama3.1:8b | 2 | 7.41% | 28 | 40.58% | 173 | 55.45% | 322 | 75.06% | 378 | 81.64% | |
| Llama3.1:70b | 7 | 25.93% | 31 | 44.93% | 228 | 73.08% | 383 | 89.28% | 439 | 94.82% | |
| GPT-4o-mini | 4 | 14.81% | 43 | 62.32% | 272 | 87.18% | 399 | 93.01% | 451 | 97.41% | |
| GPT-4o | 6 | 22.22% | 53 | 76.81% | 296 | 94.87% | 421 | 98.14% | 459 | 99.14% | |
| ThaliaType | Total FQNs | 994 | 597 | 439 | 324 | 331 | |||||
| TP | Recall | TP | Recall | TP | Recall | TP | Recall | TP | Recall | ||
| SnR | 735 | 73.94% | 533 | 89.28% | 407 | 92.71% | 300 | 92.59% | 292 | 88.22% | |
| Llama3.1:8b | 19 | 1.91% | 42 | 7.04% | 64 | 14.58% | 138 | 42.59% | 258 | 77.95% | |
| Llama3.1:70b | 18 | 1.81% | 80 | 13.40% | 144 | 32.80% | 182 | 56.17% | 270 | 81.57% | |
| GPT-4o-mini | 56 | 5.63% | 161 | 26.97% | 227 | 51.71% | 269 | 83.02% | 300 | 90.63% | |
| GPT-4o | 103 | 10.36% | 233 | 39.03% | 256 | 58.31% | 288 | 88.89% | 316 | 95.47% | |
4.2.1. Impact of Document Frequency on LLM Type Inference
To better understand why there is a large disparity between StatType-SO and ThaliaType, we investigated how the document frequency of a type impacts the performance of LLMs. Document frequency measures the number of source files that reference a specific type. Since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024), it is plausible that their performance may vary depending on the prevalence of a type within a LLMs’ training data. Understanding this bias helps ensure that LLMs deliver similar performance across all potential types used in the real world. To quantify a type’s document frequency, we analyzed source files from the Boa (Dyer et al., 2013a, b) GitHub dataset. Specifically, the document frequency is calculated as follows for a type .
Based on the document frequency, we categorized the FQNs in the StatType-SO and ThaliaType datasets into distinct frequency ranges. For each group, we calculated the recall to investigate how a type’s real-world frequency impacts LLM performance. This categorization and analysis provide insights into the extent to which LLMs are biased toward more frequently encountered types.
Results. LLMs performed poorly on less frequently used FQNs and worse on the unseen dataset ThaliaType than on StatType-SO, even for frequently used FQNs. To analyze how performance varies across FQNs with differing levels of usage frequencies, we grouped FQNs into frequency ranges by orders of magnitude (0-100, 100-1,000, 1,000-10,000, 10,000-100,000, and over 100,000 GitHub files). For the least frequently used FQNs in ThaliaType, GPT-4o only achieved a recall of 22.22% on StatType-SO and 10.36% on ThaliaType. In contrast, SnR demonstrated consistent performance regardless of a type’s popularity, achieving recalls of 100.00% and 73.94% on StatType-SO and ThaliaType respectively. On ThaliaType, LLMs only outperformed SnR on the most frequently used FQNs, specifically those appearing in over 100,000 GitHub source files. This disparity may arise from LLMs preferring more frequent FQN for a given simple name regardless of the context in the code snippet. For example, for a type with the simple name View, if android.view.View is the most frequent type in the ¿=1e5 category and javax.swing.text.View belongs to a less frequent category, then always recommending android.view.View will result in a high recall in the ¿=1e5 category but a lower recall in other categories.
5. RQ2: To what extent do LLMs understand the semantics of code snippets?
The performance decline on unseen data observed in RQ1 highlights the need to investigate how LLMs handle type inference. To explore whether this decline stems from a lack of semantic understanding in LLMs, we conducted an experiment using semantic-preserving transformations that alter syntax without changing the semantics. Our hypothesis is that if LLMs possess a deep understanding of the code snippets’ semantics, their performance should remain stable despite such transformations. Conversely, significant performance degradation would suggest that LLMs rely heavily on superficial syntactic patterns rather than a genuine semantic comprehension.
5.1. Method
The transformations, detailed in Algorithm 1, explore the extent to which syntactic changes affect LLMs’ semantic understanding of code. Three specific transformations were employed to assess the impact of modifications to identifier names (section 5.1.2), code structures (section 5.1.3), and Java keywords in comments (section 5.1.4). These were selected to highlight potential limitations in LLMs’ semantic understanding, though future research could explore additional transformations. Their impact was then evaluated using the methodology outlined in section 4. The Wilcoxon signed-rank test (Wilcoxon, 1992) was used to determine whether the transformations caused significant changes in precision, recall, and F1 scores.
5.1.1. Background: Java Grammar
The Java grammar depicted in Figure 5 is used to illustrate the key transformations on the Java code snippet. Each code snippet contains one or more classes along with a set of import statements. Classes may contain fields, methods, and blocks, with methods comprising blocks of statements and expressions. For simplicity, modifiers (e.g., public, static) were excluded, as they remain unchanged in the transformations. Other Java statements and expressions were similarly omitted as they follow the same transformation principles.
5.1.2. Identifier Renaming
This transformation investigates whether LLMs rely on specific identifier names for type inference. The RenameIdentifier function in Algorithm 1 details the process, which systematically renames identifiers such as variables, methods, classes, and packages in the code snippet. First, all variable declarations are identified using the FindVariableDeclaration function (Algorithm 1), which extracts statements matching the pattern . The algorithm then traverses each declaration, generating a unique, random three-word name for each variable using the GetRandomName function (Algorithm 1), from a predefined word list from prior work (Sotiropoulos et al., 2024) to ensure uniqueness. The algorithm replaces every occurrence of the variable name in the code (extracted using the getName function in Algorithm 1).
Next, the algorithm processes method declarations. To avoid renaming references that could alter semantics, all the method call expressions (s with the form ) that have the first optional part () is potentially external, thus added to a skip_list (Algorithm 1-1) to be filtered out (Algorithm 1). Additionally, overridden methods annotated with @Override are skipped, as their names must match those in the superclass (Algorithm 1). Lastly, the algorithm renames classes and packages (lines 1-1), following the same procedure used for variables. The details are omitted here for brevity.
| Before | After |
| p | GrouseScabsShelley |
| C | TashaMonroviaTimbers |
| a | RowsGarlickyThump |
| n() | ListerineStupefiesFetlock() |
| n | CrackerSherbertsPlod |
Figure 6 provides an example of this transformation, showing how variables, methods, classes, and packages are renamed. Method m which likely overrides the method m on type a is preserved out of an abundance of caution to preserve semantic consistency. Furthermore, it should be noted that different types of names are renamed separately. For example, the method name n is renamed to a different name than the variable name n.
5.1.3. Code Lowering
The code lowering transformation examines the impact of structural changes on LLMs’ type inference. The algorithm first extracts and traverses all the expressions (Algorithm 1) in the code snippet. For each expression, if its parent is not an expression statement (s with the form ) and it is not a loop condition (lines 1-1), the algorithm creates a new variable with a random name, assigns the expression to the variable, and inserts this assignment statement before the expression (lines 1-1). Next, the original expression is replaced with the newly created variable assigned with the value of the original expression (Algorithm 1). Such transformations modify the code structure while preserving the semantics of the code snippet. The expressions in expression statements must be skipped because transforming them may result in invalid code. For example, an expression in an expression statement can be a method call with no return value (i.e., with a void return type). In such a case, the created assignment statement would be invalid. Conditions in loops cannot be transformed either since such transformations could change the semantics of the code.
In addition to expressions, code lowering also applies to fields with an initializer. For each field in the code snippet (Algorithm 1), if it has an initializer (Algorithm 1), the algorithm splits it into a declaration and a block initializing the field (lines 1-1).
Figure 7 demonstrates this transformation. In this example, the field C c = null is split into a declaration and an initializer block with c = null. Additionally, the expression in the if statement is replaced with a newly created variable, and an assignment statement is inserted before the expression’s usage to assign its original value to the variable.
5.1.4. Adding Keyword Comments
The keyword comment transformation evaluates whether comments containing Java keywords distract LLMs. AddKeyword function in Algorithm 1 appends a single Java keyword (Gosling et al., 2023) as a line comment at the end of each line in the code snippet. An example of the transformed snippet is presented in Figure 8. These comments do not alter the semantics of the code but provide additional syntactic noise.
5.2. Results
| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| StatType-SO | 95.50% | 91.46% | 93.44% | 76.92% | 69.46% | 73.00% | 86.08% | 83.69% | 84.87% | 86.34% | 89.92% | 88.09% | 95.66% | 95.00% | 95.33% |
| Id Renaming | 95.50% | 91.46% | 93.44% | 65.28% | 56.69% | 60.68% | 85.34% | 76.15% | 80.49% | 89.63% | 88.46% | 89.04% | 97.03% | 95.62% | 96.32% |
| Code Lowering | 95.43% | 91.62% | 93.49% | 69.43% | 65.15% | 67.22% | 83.81% | 80.46% | 82.10% | 86.14% | 88.92% | 87.51% | 94.82% | 95.69% | 95.25% |
| Comment Adding | 95.50% | 91.46% | 93.44% | 78.65% | 67.46% | 72.63% | 85.39% | 80.92% | 83.10% | 87.90% | 89.38% | 88.63% | 96.06% | 95.77% | 95.92% |
| All | 95.43% | 91.62% | 93.49% | 55.83% | 47.15% | 51.13% | 77.93% | 70.08% | 73.80% | 84.60% | 82.85% | 83.72% | 93.29% | 94.15% | 93.72% |
| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| ThaliaType | 84.15% | 84.43% | 84.29% | 31.27% | 19.40% | 23.95% | 61.58% | 25.85% | 36.41% | 66.64% | 37.73% | 48.18% | 54.74% | 44.54% | 49.12% |
| Id Renaming | 84.15% | 84.43% | 84.29% | 31.36% | 19.11% | 23.74% | 57.08% | 24.47% | 34.25% | 52.47% | 38.44% | 44.37% | 43.88% | 45.77% | 44.80% |
| Code Lowering | 84.14% | 84.39% | 84.27% | 27.06% | 18.21% | 21.77% | 60.16% | 25.25% | 35.57% | 58.04% | 36.01% | 44.45% | 43.45% | 44.43% | 43.93% |
| Comment Adding | 84.15% | 84.43% | 84.29% | 25.00% | 19.22% | 21.73% | 59.56% | 25.07% | 35.28% | 64.27% | 36.57% | 46.62% | 32.26% | 44.80% | 37.51% |
| All | 84.14% | 84.39% | 84.27% | 27.57% | 17.77% | 21.61% | 45.92% | 23.87% | 31.41% | 47.50% | 36.87% | 41.52% | 56.30% | 45.29% | 50.20% |
The effects of applying transformations to the StatType-SO and ThaliaType datasets are shown in Tables 3(a) and 3(b), respectively. These transformations significantly impacted the performance of all LLMs, resulting in significant decreases in both precision and recall.
When all three transformations were applied simultaneously, the decrease in performance for LLMs was amplified. The F1-score dropped significantly () for all LLMs on StatType-SO (e.g., 1.7% decrease on GPT-4o, and 30.0% decrease on Llama3.1:8b). Precision dropped significantly for all LLMs across both the StatType-SO and ThaliaType datasets, while recall also declined on both datasets for nearly all models. Comparing the F1-score changes revealed that LLMs were more affected by the transformations in StatType-SO than in ThaliaType. Unlike ThaliaType, which consists entirely of unseen code snippets, the more substantial decline in StatType-SO indicates potential data leakage, where LLMs rely on memorized patterns rather than demonstrating a deep semantic understanding of the code.
In contrast, SnR, which analyzes the semantic meaning behind the code snippets, was unaffected by these transformations, experiencing only minor, statistically insignificant variations under the code lowering transformation. These variations occurred when multiple types satisfied all constraints, with the order of type variables serving as a tie-breaker.
6. RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
Building upon the findings of RQ2, which demonstrated the sensitivity of LLMs to subtle syntactic variations, we investigate the fundamental sources of LLM’ type inference performance. Specifically, we aim to identify the minimal set of syntax elements sufficient for LLMs to perform type inference.
6.1. Method
We applied delta debugging to systematically minimize the input code snippet while preserving consistent inference results. This approach enables us to isolate the exact syntactic elements that are crucial for LLMs’ type inference. The intuition is that tokens conveying more semantic information and critical for the type inference task, such as simple names, should be preserved, whereas tokens carrying less semantic information, such as whitespaces, can be removed. Identifying these crucial syntactic elements provides valuable insights into the factors that influence LLM performance in type inference and can inform the development of more robust and efficient type inference systems.
6.1.1. Background: Delta Debugging
Delta debugging (Zeller and Hildebrandt, 2002), specifically the ddmin algorithm, is a widely used technique in software debugging to minimize a given input while preserving a certain property. In our study, the property we intend to preserve is that the minimized code snippet contains sufficient information for LLMs to infer the same import statements as the original code snippet. The minimized code snippet preserves this property, representing the elements in the code snippet sufficient for LLMs to conduct type inference.
In this research question, in addition to ddmin, we also used Perses (Sun et al., 2018) to minimize code snippets due to its high efficiency. Notably, Perses achieves this efficiency by ensuring the transformed program remains syntactically valid. However, since LLMs can interpret syntactically invalid programs, we apply ddmin to further minimize the output from Perses. Unlike in previous RQs, the purpose of the minimization here is to identify the exact syntax elements required for LLMs’ type inference, meaning that syntactical validity does not need to be preserved. In the end, the minimized program is 1-minimal—meaning that the removal of any single token will cause the type inference tool to fail in inferring the same import statements.
6.1.2. Code Snippet Minimization
Algorithm 2 outlines the core conceptual workflow of code snippet minimization. The minimization process follows a trial-and-error approach as defined by the Minimize function. During this process, a minimal snippet min is maintained and initially set to the original snippet (Algorithm 2). Given the current minimal snippet min, a delta debugging algorithm generates a set of smaller snippets that may retain the property, i.e., the LLM infers the same import statements as the original ones for the snippet (Algorithm 2). The primary differences among different delta debugging algorithms lay in the strategy of generating smaller snippets. For brevity, we abstract the distinctions between the delta debuggers, (i.e., ddmin and Perses) using the ComputeSmallerSnippets function. If any of the smaller snippets preserve the property, the current minimal snippet min is updated, and the process continues (lines 2-2). Otherwise, if none of the smaller snippets preserve the property, indicating that no further progress can be made, the minimization terminates (lines 2-2). In our experiment, to minimize a code snippet, the import statement inferred by the LLM is first extracted (Algorithm 2), after which, the algorithm applies snippet minimization with both Perses and ddmin to derive the final minimal snippet (lines 2-2). Figure 9 illustrates this process using a code snippet from StatType-SO (Figure 9(a)). In this example, GPT-4o-mini initially inferred four types: Session, Transaction, SessionFactory, and Configuration (Figure 9(b)). After applying Algorithm 2, the snippet was minimized to a single line ibernate (Figure 9(c)).
The minimization process is computationally expensive, particularly the repeated type inference performed for each smaller variant (Algorithm 2). The runtime for each type inference operation varies depending on the model’s size and complexity. With tens of thousands of inference calls required, due to cost reasons, we focused on GPT-4o-mini only for this research question.
6.2. Result
The minimization process eliminated a greater portion of tokens from code snippets in StatType-SO compared to ThaliaType, as summarized in Figure 10. On average, only 6.37% of the tokens were retained for each code snippet in StatType-SO, whereas 14.73% of the tokens were preserved for each code snippet in ThaliaType, more than double. While the original datasets contained 95803 and 91166 tokens for StatType-SO and ThaliaType respectively, the minimized versions only contained 4619 and 16317 tokens (Table 4). The higher removal ratio indicates that many of the tokens in the original code snippet are not required to achieve the high type inference performance seen in StatType-SO.
| StatType-SO | ThaliaType | |||||
| Category | Ratio | # in Original | # in Minimized | Ratio | # in Original | # in Minimized |
| Average | 4.82% | 95803 | 4619 | 17.90% | 91166 | 16317 |
| Control and Structural | 3.17% | 36791 | 1165 | 14.39% | 37864 | 5450 |
| Inferred Simple Name | 16.88% | 4954 | 836 | 53.33% | 4478 | 2388 |
| Whitespace | 6.18% | 29388 | 1816 | 20.95% | 29144 | 6105 |
| Other | 3.25% | 24670 | 802 | 12.06% | 19680 | 2374 |
To analyze how different types of tokens are affected by the minimization process, tokens are categorized into four groups, control and structural, whitespace, inferred simple name, and other. The control and structural category includes symbols (e.g., parenthesis, semicolon), operators (e.g. plus, minus, equals), and keywords (e.g. if, while), as defined by the Java Language Specification (JLS, 2021). The whitespace category consists of tokens composed entirely of whitespace characters (e.g., space, new line, tab). The inferred simple name category includes simple names of the types inferred by the LLM. If the simple name is not inferred, it falls into the other category, which also includes tokens such as integer or floating-point literals. As shown in Table 4 control and structural tokens, along with those in the other category, are frequently removed during minimization. In contrast, whitespace and inferred simple name tokens are often retained.
We further analyzed the inferred simple name category to determine whether simple names were consistently preserved in minimized code snippets. Retaining these names is crucial for the task of inferring FQNs of types, since such types are often referenced by their simple names. However, as shown in Figure 11, these simple names are frequently absent in minimized versions. To quantify this effect, we calculated the percentage decrease in the unique simple names after minimization. To account for simple names that were not inferred, we only included those for which LLMs provided an FQN. For example, while the code snippet in Figure 9(a) contained all four of the inferred type’s simple names (Figure 9(b)), none were preserved in the minimized version (Figure 9(c)), representing a 100% decrease. On average, the number of unique simple names in StatType-SO decreased by 41.49% , suggesting that inferring the types in StatType-SO often does not require simple names. In contrast, ThaliaType exhibited only a 9.12% decrease.
These findings suggest that LLMs may rely more on direct recollection of their training data than truly interpreting the semantic information in code snippets in StatType-SO. Since type inference inherently requires the simple names to resolve FQNs, the ability of LLMs to infer types without these simple names indicates that they may not be solving the task as intended. Instead, their high performance may stem from exploiting superficial patterns in code snippets. This highlights a weakness in the StatType-SO dataset: its potential predictability and insufficient penalization of overfitting. Consequently, LLM performance on StatType-SO may not generalize effectively to unseen code snippets. To ensure a more robust and generalizable assessment of LLM performance, future work should focus on evaluating models using datasets composed entirely of unseen code snippets.
7. Threats to Validity
There are three main threats to validity. First, variations in the document frequency of FQNs between ThaliaType and StatType-SO datasets could influence LLM type inference performance, since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024). LLM’s performance may vary depending on the prevalence of a type within LLM’s training data. To address this, we analyzed LLM performance across different document frequency levels. Despite these efforts, LLMs consistently exhibited lower performance on unseen code snippets from ThaliaType compared to StatType-SO.
Second, our findings may not generalize to other LLMs. We mitigated this by following best practices (Sallou et al., 2024) and evaluated a diverse set of models, including both open weight (Llama3.1:8b, Llama3.1:70b) and the state-of-the-art closed models (GPT-4o, GPT-4o-mini) of varying sizes, providing a representative sample of current LLM capabilities. However, due to the high cost of performing minimization, we were only able to evaluate GPT-4o-mini for RQ3. Nevertheless, since all current LLMs are based on transformers, their underlying type inference mechanisms operate in a similar manner. Thus our findings in RQ3 remain generalizable—specifically, that the StatType-SO dataset is insufficient for evaluating type inference.
Third, prompt engineering may enable LLM to achieve better performance for type inference on code snippets. To mitigate this threat, we followed best practices to design a simple but effective prompt (LLM, 2024) which achieved high performance on the StatType-SO dataset. Regardless, data leakage remains a fundamental issue that can influence LLM evaluations regardless of prompt quality. Our evaluation highlights how data leakage may have impacted prior assessments of LLM type inference performance.
8. Related Work
In this section, we discuss related work in the fields of type inference for code snippets and LLM data leakage.
8.1. Type Inference for Code Snippets
Existing type inference approaches for code snippets mainly fall into two categories, constraint-based approaches and machine-learning-based approaches.
Constraint-Based Approaches. The general idea of constraint-based approaches is to first analyze the code snippet and derive a collection of constraints on the types that need to be inferred. With the derived constraints, a set of APIs that satisfies all the constraints can be obtained by performing a constraint-solving algorithm. Baker is the first work that applied constraint-based type inference for code snippets (Subramanian et al., 2014). A subsequent study, SnR, outperformed Baker by improving the handling of parameterized types and leveraged Datalog to efficiently solve the derived constraints against a knowledge base of known types (Dong et al., 2022).
Machine-Learning-Based Approaches. Machine-learning-based approaches typically leverage a model trained with a large set of programs from open-sourced projects. Phan et al. (2018) proposed StatType, which learns the FQNs that often co-occur from a large corpus. With such knowledge, StatType can derive the FQN for an API based on the neighboring API names. A subsequent work conducted by Saifullah et al. (2020) made improvements by leveraging both local and global contexts. Huang et al. (2023a) leveraged pre-trained CodeBert (Feng et al., 2020), a transformer-based masked language model to predict FQNs in the code snippet. Compared to LLMs evaluated in this study, CodeBert is much smaller at 125 million parameters compared to even Llama3.1:8b’s 8 billion parameters. Kabir et al. (2024) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt. However, since our work focuses on mitigating data leakage, which fundamentally undermines inference reliability regardless of the prompting strategy used, we used a single prompt.
8.2. Type Inference for Dynamically Typed Languages
Type inference for dynamically typed languages faces extra challenges as there can be variables with insufficient static type constraints in programs written in these languages, which static type inference cannot soundly handle. Many existing studies resort to deep learning-based approaches to overcome this challenge. For example, Hellendoorn et al. (2018) propose DeepTyper, a deep learning model that is trained to provide type suggestions given certain contexts and relations. Pradel et al. (2020) propose TypeWriter, which uses a deep learning-based approach to predict the types and then utilizes a search-based approach to validate the predicted type. Peng et al. (2022) propose HiTyper, which is a deep learning-based approach combining static type inference. It conducts static inference and DL-based prediction iteratively to construct a type dependency graph, which records type dependency information among variables. This paper focuses on the performance of LLMs in type inference for Java code snippets, leaving the investigation of type inference for dynamically typed languages to future work.
8.3. LLM Data Leakage
Data leakage is not limited to type inference but also affects other software engineering tasks, such as automated program repair using LLMs (Xia et al., 2023; Ouyang et al., 2024). Xia et al. (2023) discovered that 15% of the bug-fixing patches for automated program repair generated by earlier LLMs (i.e., CodeT5, GPT-Neo, GPT-J, and GPT-NeoX) were already present in the training data. Sainz et al. examined several academic datasets and reported that most of them were either used as training data for ChatGPT or likely exposed to it (Sainz et al., 2023). A subsequent work done by Golchin and Surdeanu (2023) proposes a more advanced detection technique and detects the presence of test data of several datasets in the training data of LLMs. As LLM training datasets grow larger and increasingly opaque, detecting instances of data leakage during evaluation becomes progressively more challenging.
The data leakage problem remains an open challenge for evaluating LLMs. Some studies have been conducted in other fields (Magar and Schwartz, 2022; Balloccu et al., 2024; Mirzadeh et al., 2024; Cao et al., 2024) regarding data leakage and LLM evaluation. Mirzadeh et al. (2024) showed that LLMs have significant limitations when conducting genuine mathematical reasoning which suggests that LLMs are conducting sophisticated pattern matching rather than true logical reasoning. A recent study (Cao et al., 2024) demonstrated that the performance of LLMs is far behind undergraduate students on the proposed project-level Java benchmark that exercises OOP features. Our evaluation on type inference rather than math questions or code generation also showed a significant drop in performance with generated code snippets which suggests that understanding the type system in code is challenging for LLMs. So far, there is no conclusive evidence showing that LLMs are conducting true type inference on code snippets. Small changes to the inputs can drastically alter model outputs (Shi et al., 2023; Mirzadeh et al., 2024; Jiang et al., 2024). While some guidelines have been raised for software engineering research using LLMs (Ozkaya, 2023; Sallou et al., 2024), protecting against data leakage remains an open challenge. We hope our work will shed further light on the data leakage issue in software engineering.
9. Conclusion
This paper conducted a comprehensive assessment of LLMs’ type inference capabilities on Java code snippets. Our three-pronged evaluation mitigated potential data leakage issues and identified possible limitations of LLMs. First, we create a new dataset named ThaliaType. By evaluating the performance of LLMs on ThaliaType and StatType-SO, we conclude that LLM type inference performance may not generalize well to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall. In addition, LLM type inference performance diminishes when inferring less commonly used FQNs, potentially introducing bias and presenting opportunities for future research to enhance performance. Second, we designed and applied three semantic-preserving code transformations to code snippets in ThaliaType and StatType-SO to investigate LLMs’ semantic understanding. Through evaluating LLMs with these transformed code snippets, we find that the precision and recall of LLMs in type inference depend on the exact content of the code snippet, even when the semantic meaning of the code remains unchanged. This suggests that LLMs may rely heavily on superficial patterns in the code rather than truly understanding its semantics, harming LLMs’ performances’ generaliability. Third, we apply delta debugging to analyze the minimal syntactic elements sufficient for LLMs to infer types. We find that LLMs can still correctly infer FQNs even when the types were absent from the snippets, that StatType-SO may allow LLMs to rely on superficial patterns rather than deep code understanding. Our experiments suggest that LLMs’ strong performance in prior evaluations was likely due to data leakage rather than a genuine understanding and analysis of the semantics in code snippets. Our findings suggest that future evaluations of LLMs should utilize a generated dataset, carefully designed to assess the true capabilities of LLMs for type inference tasks. All of our code and dataset are available at https://github.com/uw-pluverse/thalia-type.

### v2

@@
-Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets
+Unmasking the Genuine Type Inference Capabilities of LLMs for Java Code Snippets
 Abstract.
-Type inference is a crucial task for reusing online code snippets, often found on platforms like StackOverflow, which frequently lack essential type information such as fully qualified names (FQNs) and required libraries.
+Type inference is a crucial task for reusing online code snippets.
+Although code snippets are prevalently shared on platforms like StackOverflow, they frequently lack essential type information, such as fully qualified names (FQNs) and required libraries.
 Recent studies have leveraged Large Language Models (LLMs) to perform type inference for such code snippets, demonstrating promising performance.
-However, these results are potentially affected by a data leakage issue, as the benchmark suite (StatType-SO) used for evaluation has been publicly available on GitHub since 2017, with the full suite available since 2023.
-Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data.
-To comprehensively assess the type inference capabilities of LLMs on Java code snippets and identify potential limitations of LLMs, we conducted a three-pronged evaluation.
-First, utilizing Thalia, a program synthesis technique, we created ThaliaType—a new, previously unreleased dataset designed for type inference evaluation.
-ThaliaType comprises 300 unseen code snippets, generated in a manner consistent with the setting of StatType-SO.
-Our results revealed that LLMs exhibit a significant drop in performance when generalizing to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall.
-To further investigate the limitations of LLMs in understanding the semantic meaning of the code, we developed semantic-preserving code transformations.
-Our analysis showed that LLMs performed significantly worse on code snippets that are syntactically different but semantically equivalent.
-Finally, we employed delta debugging to distill the minimal syntax elements in snippets sufficient for inference by LLMs.
-We observe that while the type inference task primarily involves inferring FQNs for types used in code snippets, LLMs could still correctly infer FQNs even when the types were absent from the snippets.
-This suggests that LLMs rely more on re-using knowledge from training data instead of thoroughly analyzing the snippets.
-Our experiments suggest that the strong performance of LLMs in prior evaluations was likely influenced by data leakage in the benchmarks, rather than a genuine understanding of the semantics of code snippets.
-Our findings highlight the crucial need for carefully designed and rigorously evaluated benchmarks to accurately assess the true capabilities of LLMs for type inference tasks.
-To ensure reliable and rigorous evaluations of LLMs’ type inference capabilities, future research should prioritize the use of unseen benchmarks that are explicitly excluded from the LLMs’ training data.
+However, these evaluations may suffer from data leakage, as the benchmark suite, StatType-SO, used for evaluation has been publicly available on GitHub since 2017.
+Consequently, it remains uncertain whether the strong performance of LLMs reflects genuine semantic understanding of code or is due to the ground truth being included in the training set.
+This paper strives to comprehensively evaluate the genuine type inference capabilities of LLMs on Java code snippets and identify potential limitations of LLMs.
+First, we created ThaliaType, a new, previously unreleased benchmark suite designed for type inference evaluation.
+Second, using the StarCoder2 LLM as a baseline, we uncovered data leakage from StatType-SO in StarCoder2’s open-source training set and observed that other state-of-the-art LLMs exhibit similar performance drops when evaluated on ThaliaType, with precision decreasing by up to 59% and recall by up to 72%.
+Finally, we developed semantic-preserving code transformations to further investigate the capabilities of LLMs in understanding the execution semantics of code snippets.
+Our results showed that the performance of LLMs on StatType-SO is far less robust to these transformations than on ThaliaType, suggesting that the performance on StatType-SO may be biased by data leakage and have limited generalizability.
+These findings highlight the importance of carefully designed, leakage-free benchmarks for evaluating LLMs on type inference tasks.
+We recommend future studies adopt ThaliaType to ensure rigorous and reliable assessments of LLMs’ genuine type inference capabilities.
 1. Introduction
@@
 For instance, if a code snippet references a type with the simple name Logger, the developer must determine whether the appropriate import is java.util.logging.Logger from Java’s utilities library or org.slf4j.Logger from another logging library, depending on the methods invoked on the objects of Logger in the code snippet.
-Type inference for code snippets can assist developers in recovering missing type information needed to reuse these code snippets (Dong et al., 2022; Subramanian et al., 2014; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023b; Kabir et al., 2024).
-Existing approaches to type inference for Java code snippets generally fall into two categories: constraint-based and machine learning (ML)-based techniques.
+Type inference for code snippets can assist developers in recovering missing type information needed to reuse these code snippets (Dong et al., 2022; Subramanian et al., 2014; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023b; Kabir et al., 2025).
+For example, current approaches to type inference for Java code snippets generally fall into two categories: constraint-based and machine learning (ML)-based techniques.
 Constraint-based methods build a knowledge base of libraries and the types contained within (Dong et al., 2022; Subramanian et al., 2014).
 The state-of-the-art technique SnR (Dong et al., 2022) leverages this knowledge base and uses constraints extracted from the code snippet to identify the precise types being used.
-In contrast, ML-based approaches (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024) learn from prior examples of type usage in existing code to perform type inference.
-ML-based techniques often produce incorrect inferences, due to their limited understanding of code structure and the rules governing Java’s type system.
+ML-based approaches (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025) learn from prior examples of type usage in existing code to perform type inference.
+They often produce incorrect inferences due to their limited understanding of code structure and the rules governing Java’s type system.
 Recent advancements in large language models (LLMs), such as OpenAI’s proprietary GPT family and Meta’s open-weight Llama models, offer new possibilities for overcoming these limitations.
 By leveraging training on diverse and extensive datasets, LLMs have shown potential in performing various software engineering tasks (Hou et al., 2024).
-Prior studies (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) suggest that LLMs achieved performance comparable to that of the state-of-the-art, SnR.
-However, the evaluation of the prior techniques (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) is potentially affected by data leakage (Inan et al., 2021) where the training dataset includes the benchmark dataset, and data leakage allows LLMs to regurgitate the training data rather than conducting type inference.
-The benchmark suite, StatType-SO, used for evaluating type inference techniques has been publicly available on GitHub since 2017 (Sta, 2017), with the full suite available since 2023 (Sta, 2023).
+Prior studies (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) suggest that LLMs achieved performance comparable to that of the state-of-the-art SnR.
+| Benchmark Suite | First Available | #Code Snippets |
+| StatType-SO | 2017 (see Table 2) | 267 |
+| ThaliaType | Always New | 300 |
+| Model | Version | Knowledge Cutoff | Data Leakage |
+| StarCoder2:15b | instruct-v0.1-q4_0 | September 2023 (Lozhkov et al., 2024) | ✓ |
+| GPT-4o | 2024-08-06 | October 2023 (OPE, 2024a) | ? |
+| GPT-4o-mini | 2024-07-18 | October 2023 (OPE, 2024a) | ? |
+| Llama3.1:8b | instruct_q4_0 | December 2023 (Meta, 2024) | ? |
+| Llama3.1:70b | instruct_q4_0 | December 2023 (Meta, 2024) | ? |
+However, the evaluation of these prior studies (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) is potentially affected by data leakage (Inan et al., 2021) where the training dataset includes the benchmark dataset together with the ground truth, and data leakage allows LLMs to regurgitate the training data rather than conducting type inference.
+The benchmark suite StatType-SO, used for evaluating type inference techniques has been fully, publicly available on GitHub since 2017 (Sta, 2017).
 As shown in Table 1, recent state-of-the-art LLMs have knowledge cutoff dates in late 2023, creating the potential for leakage.
-Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data.
-Since the exact training data for LLMs are kept confidential, detecting such leakage is difficult (Carlini et al., 2021; Nasr et al., 2023).
+Thus, it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data.
 While recent work has called attention to this data leakage problem for software engineering research (Ozkaya, 2023; Sallou et al., 2024), no study has thoroughly, empirically evaluated LLMs’ performance on type inference on code snippets.
-In this paper, we thoroughly evaluate LLMs’ type inference performance on Java code snippets to better understand the strengths and limitations of LLMs on type inference tasks, and uncover potential limitations of LLMs.
-We organized our inquiry into the following three research questions.
+Because prior work relied on GPT models where the exact training data is kept confidential, detecting such leakage is difficult (Carlini et al., 2021; Nasr et al., 2023).
+To address this limitation, we first investigated The Stack v2 (BigCode, 2024) (simply referred to as StackV2), an open-source dataset used to train the StarCoder2 model.
+Our exploration revealed that all code snippets from StatType-SO with ground truth were present in StackV2.
+This means that StarCoder2 was trained using code snippets along with their expected import statements from StatType-SO.
+Thus, StarCoder2’s performance is affected by data leakage, and can potentially recall the code snippets and their expected import statements from training rather than performing type inference by analyzing the semantics of the code snippets.
+To better understand the behavior of LLMs, we compare StarCoder2’s performance against Llama and GPT models.
+We organized our inquiry into the following two research questions as illustrated in Figure 1.
+In this paper, we thoroughly evaluate LLMs’ type inference performance on Java code snippets, focusing on their ability to leverage type information, the impact of data leakage, and their generalizability to unseen code.
 RQ1: How well do LLMs perform type inference on unseen code snippets?
-To address data leakage, we generated a new dataset named ThaliaType using Thalia, a program synthesis technique originally designed for testing static type systems.
-ThaliaType comprises 300 code snippets guaranteed to be unseen by the evaluated LLMs, thereby avoiding data leakage.
-Furthermore, we used the same set of libraries as used in StatType-SO to generate ThaliaType, to (1) enable a direct comparison of type inference performance between the two datasets, and (2) ensure the evaluated LLMs were consistently and sufficiently trained on the same set of types used in both datasets.
-By evaluating LLMs on ThaliaType and contrasting those results with those from StatType-SO, we can better understand the impact of data leakage on performance.
-RQ2: To what extent do LLMs understand the semantics of code snippets?
-To investigate whether LLMs truly understand the semantics of code snippets during type inference, we designed semantic-preserving code transformations to create syntactically different versions of the input code (section 5).
+To address data leakage, and investigate LLMs’ genuine ability to perform type inference using the semantic information present in code snippets, we generated a new benchmark suite named ThaliaType using Thalia, a program synthesis technique originally designed for testing static type systems (Sotiropoulos et al., 2024).
+ThaliaType comprises 300 code snippets that are guaranteed to be unseen by the evaluated LLMs, thereby avoiding data leakage.
+Unlike StatType-SO, which was published in 2017 (Table 1(a)), new code snippets can always be generated using our replication package, enabling future evaluations without data leakage.
+To generate ThaliaType, we used the same set of libraries used in StatType-SO to (1) enable a direct comparison of type inference performance between the two benchmark suites and (2) ensure the evaluated LLMs were consistently and sufficiently exposed to the types used in both benchmark suites during training.
+By evaluating LLMs on ThaliaType and contrasting the results with those from StatType-SO, we can gain a better understanding of the impact of data leakage on performance.
+RQ2: To what extent do LLMs understand the execution semantics of code snippets?
+To investigate whether LLMs truly understand the execution semantics of code snippets during type inference, we designed semantic-preserving code transformations to create syntactically different versions of the input code (section 6).
 These transformations were intended to hinder LLMs’ reliance on memorized training data by making direct recall more challenging.
-By observing how LLMs’ performance changes on these semantically equivalent but syntactically distinct versions, we can gain insights into their ability to grasp the underlying semantic meaning of the code.
-RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
-Lastly, we sought to distill the minimal syntax elements sufficient for LLMs’ inference through delta debugging (Zeller and Hildebrandt, 2002) (section 6).
-By minimizing the input code to the LLMs while maintaining the LLMs’ inference results, we assess whether LLMs rely on superficial patterns from training data rather than genuinely analyzing code semantics.
-If LLMs are truly performing type inference rather than leveraging memorized patterns, essential elements such as simple names should be retained in the minimized snippets.
-Following the recommendation from prior software engineering research (Sallou et al., 2024), we evaluate LLM’s performance using two open-source models Llama3.1:8b and Llama3.1:70b, alongside state-of-the-art closed-source models GPT-4o and GPT-4o-mini (see Table 1(b)).
-As a baseline, we compare these models against SnR (Dong et al., 2022), the state-of-the-art constraint-based type inference technique.
-This evaluation aims to provide deeper insights into the potential limitations of using LLMs and to guide future research in evaluating LLMs’ performance on other software engineering tasks.
-In RQ1, we found LLMs are highly likely to be affected by the data leakage issue and are prone to making incorrect inferences on unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall for Llama3.1:8b.
-On ThaliaType, even the best LLM GPT-4o only obtained a precision of 54.74% and a recall of 44.54% .
-On the other hand, SnR not affected by the data leakage issue, obtained a precision of 84.15% and a recall of 84.43% on ThaliaType.
-This stark discrepancy suggests LLMs are affected by data leakage, thus prompting further investigation into their type inference performance.
-In RQ2, we found that LLMs performed significantly worse on transformed code snippets that were syntactically different but semantically equivalent.
-Applying all transformations caused a significant drop in F1 scores on the StatType-SO dataset across all LLMs.
-This suggests that LLMs rely heavily on the exact syntactic structure, indicating a lack of semantic understanding that limits their generalizability in type inference.
-In RQ3, delta debugging revealed that LLMs can infer the same types using only 6.37% of the tokens in StatType-SO, with the number of unique simple names decreasing by 41.49% after minimization.
-Despite the type inference task primarily involves inferring FQNs for the types in the code snippet, LLMs would still correctly infer FQNs even when the types were absent.
-This suggests that StatType-SO may enable LLMs to perform well by exploiting superficial patterns in code snippets; this performance may be due to data leakage rather than a genuine understanding of code semantics.
-Our experiments indicate that prior evaluations of LLMs were likely influenced by data leakage rather than a genuine comprehension of the semantics of the code snippets.
-To address these in the future, it is essential to evaluate LLMs on benchmarks that effectively measure their performance and are guaranteed to be excluded from the LLMs’ training data such as using Thalia to generate unseen code snippets.
+By observing how LLM’ performance changes on these semantically equivalent but syntactically distinct versions, we can gain insights into their ability to grasp the underlying semantic meaning of the code.
+Following recommendations from prior software engineering research (Sallou et al., 2024), we evaluate LLM’s performance using three open-weight models StarCoder2, Llama3.1:8b, and Llama3.1:70b, along with state-of-the-art closed-source models GPT-4o and GPT-4o-mini (see Table 1(b)).
+Among these, StarCoder2 serves as a baseline model with confirmed data leakage (see section 4).
+For additional comparison, we include SnR (Dong et al., 2022), a state-of-the-art constraint-based type inference technique, as a non-LLM baseline.
+This evaluation aims to provide deeper insights into the potential limitations of using LLMs and to guide future research evaluating LLMs’ performance on other software engineering tasks.
+Findings of RQ1.
+In RQ1, we found that LLMs are prone to making incorrect inferences on code snippets from ThaliaType, likely because these snippets are truly unseen and not affected by data leakage.
+All tested LLMs exhibited a similar decline in performance as StarCoder2:15b which experienced a 66.8% decrease in F1-score.
+For example, Llama3.1:8b and GPT-4o showed comparable declines of 67.2% and 48.5% respectively.
+On both StatType-SO and ThaliaType, StarCoder2:15b outperformed Llama3.1:8b but was outperformed by Llama3.1:70b, GPT-4o-mini, and GPT-4o.
+In contrast, SnR, which is not affected by data leakage, obtained a precision of 84.15% and a recall of 84.43% on ThaliaType.
+The consistent drop in LLMs’ performance between the baseline StarCoder2:15b and other models suggests that all tested LLMs are potentially affected by data leakage.
+In addition, although StarCoder2:15b was trained on StatType-SO snippets, it only achieved an F1-score of 81.67% , indicating that even in the presence of data leakage, models do not perfectly recall the training data.
+Findings of RQ2.
+In RQ2, LLMs showed robustness to individual transformations on both benchmark suites, and also to the combined transformation on ThaliaType code snippets.
+StarCoder2:15b showed only a 3.0% decrease in F1-score in the worst case, while Llama3.1:8b experienced a 7.9% drop.
+However, combined transformations on StatType-SO consistently led to performance degradation across all models, including a 16.8% reduction for StarCoder2 and up to 30.0% for Llama3.1:8b.
+Notably, these same combined transformations did not consistently degrade performance on ThaliaType.
+In fact, StarCoder2:15b experienced a slight improvement in F1-score.
+One possible explanation is the presence of data leakage, which may allow models to recall answers from training.
+When transformations introduce enough variation to prevent direct recall, the models are forced to rely on semantic reasoning.
+Under these conditions, their performance on StatType-SO aligned more closely with that on truly unseen data.
+Our results suggest that future evaluations of LLM-based techniques should also assess generalizability, defined as the ability for LLM-based techniques to maintain consistent performance across unseen code snippets.
+This can be achieved by incorporating both transformed code snippets and ThaliaType, rather than relying solely on StatType-SO, which may suffer from data leakage.
 Contribution. This paper makes the following contributions.
-We proposed a new dataset ThaliaType along with an approach to creating such new datasets to rigorously evaluate type inference performance of LLMs while mitigating the potential biases introduced by data leakage.
-With the help of ThaliaType, we demonstrated that prior evaluations using StatType-SO are likely subjected to the data leakage problem and thus the performance results are inflated and misleading.
-We investigated how well LLMs understand the semantics of code snippets by experimenting with whether semantic-preserving transformations on the code snippets affect the type inference performance of LLMs.
-The evaluation results on StatType-SO and ThaliaType show that the precision and recall of LLMs on type inference are dependent on the exact content in the code snippet despite preserving the semantic meaning of the code, indicating the LLMs lack a deep comprehension of the semantics and heavily rely on superficial syntactic patterns for type inference.
-We further investigated the essence of LLMs performing type inference by isolating the syntactic elements sufficient for LLMs’ using delta debugging.
-The results show that some essential simple names are not required by LLMs for type inference especially on StatType-SO, indicating LLMs may rely more on exploiting superficial patterns from their training data than on the semantic information in code snippets.
-This study provides a novel approach to rigorously assess LLMs’ comprehension of program semantics.
-To enable replication, reproduction, and further research on LLMs for code-related tasks, we have made our dataset and code publicly available at https://github.com/uw-pluverse/thalia-type.
+We introduce ThaliaType, a new benchmark suite constructed through a novel adaptation of Thalia for rigorously evaluating the type inference performance of LLMs while mitigating potential bias from data leakage.
+Using ThaliaType, we demonstrate that prior evaluations based on StatType-SO are likely affected by data leakage, leading to inflated and potentially misleading performance results.
+We evaluate whether LLMs understand the execution semantics of code snippets by applying transformations that preserve the semantics and comparing their effects.
+Our results on StatType-SO and ThaliaType show that LLMs are generally able to maintain performance across individual transformations.
+However, complex transformations on StatType-SO result in disproportionately large performance drops compared to both their effects on ThaliaType and the individual transformations.
+This lack of generalizability can be caused by data leakage where models learned the correct answer from training.
+To enable generation of new code snippets for future research on LLM type inference for code snippets, we have made our benchmark suite and code publicly available at https://github.com/uw-pluverse/thalia-type.
 2. Type Inference on Java Code Snippets
-Type inference on code snippets aims to automatically determine the precise types within the code, specifically their FQNs (Subramanian et al., 2014; Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024).
+Type inference on code snippets aims to automatically determine the precise types within the code, specifically their FQNs (Subramanian et al., 2014; Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025).
 This process is essential for effectively reusing code snippets from sources like StackOverflow, where type information is often incomplete or missing entirely.
+| Simple Name | Inferred FQN |
+| String | java.lang.String |
+| Logger | java.util.logger.Logger |
+| Simple Name | FQN | Method Name | Return Type | Argument Types |
+| Logger | java.util.logging.Logger | getLogger | java.util.logging.Logger | java.lang.String |
+| Logger | java.util.logging.Logger | info | void | java.lang.String |
+| Logger | org.slf4j.Logger | info | void | java.lang.String |
+| Logger | org.slf4j.Logger | debug | void | java.lang.String |
+| String | java.lang.String | length | int | |
 A key challenge in type inference lies in resolving ambiguities with simple names.
@@
 2.1. Machine Learning-Based Type Inference
-Machine learning-based methods use statistical models and learned representations of code to conduct type inference and resolve ambiguous simple names in code snippets (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024).
-Of these, LLMs have recently achieved state-of-the-art performance on type inference on the StatType-SO (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) dataset.
-Existing approaches typically employ a simple but effective prompt to guide LLMs in generating FQNs for types in the code snippet (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024).
+Machine learning-based methods use statistical models and learned representations of code to conduct type inference and resolve ambiguous simple names in code snippets (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025).
+Of these, LLMs have recently achieved state-of-the-art performance on type inference on the StatType-SO (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) benchmark suite.
+Existing approaches typically employ a simple but effective prompt to guide LLMs in generating FQNs for types in the code snippet (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024).
 Notably, Kabir et al.
 (2025) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt.
-This iterative process achieved a precision of 99.5% on the StatType-SO dataset.
+This iterative process achieved 99.5% precision on StatType-SO.
 In contrast, our work focuses on the issue of data leakage in type inference rather than repairing code snippets.
@@
 To ensure consistency and comparability with prior research, we adopted a single-prompt approach for type inference.
-Following established practices (Huang et al., 2023a; Kabir et al., 2024), we developed a straightforward but highly effective prompt for LLM-based type inference.
+Following established practices (Huang et al., 2023a; Kabir et al., 2025), we developed a straightforward but highly effective prompt for LLM-based type inference.
 Prompt Template.
@@
 Despite its simplicity, our prompt achieves high accuracy and aligns with best practices for LLM prompt engineering (LLM, 2024).
-Notably, our result using GPT-4o on the StatType-SO dataset represents a 7% improvement over the prior results (Kabir et al., 2024), demonstrating the robustness of our prompt while maintaining a similar level of complexity.
+Notably, our results using GPT-4o on the StatType-SO benchmark suite represent a 7% improvement over the prior results (Kabir et al., 2025), demonstrating the robustness of our prompt while maintaining a similar level of complexity.
 | System | You are a helpful programming assistant. |
@@
 SnR (Dong et al., 2022) is the state-of-the-art technique that uses constraint-based type inference (Dong et al., 2022; Subramanian et al., 2014).
-It has proven highly effective for inferring types in code snippets from the StatType-SO dataset.
+It has proven highly effective for inferring types in code snippets from StatType-SO.
 To illustrate the information SnR leverages, consider the example code in Figure 2(a).
@@
 These findings will provide deeper insights into LLMs’ type inference capabilities and highlight opportunities for future research.
-| Simple Name | Inferred FQN |
-| String | java.lang.String |
-| Logger | java.util.logger.Logger |
-| Simple Name | FQN | Method Name | Return Type | Argument Types |
-| Logger | java.util.logging.Logger | getLogger | java.util.logging.Logger | java.lang.String |
-| Logger | java.util.logging.Logger | info | void | java.lang.String |
-| Logger | org.slf4j.Logger | info | void | java.lang.String |
-| Logger | org.slf4j.Logger | debug | void | java.lang.String |
-| String | java.lang.String | length | int | |
-3. Datasets
-We utilize two datasets StatType-SO and ThaliaType for evaluating LLMs’ type inference performance on code snippets.
+3. Benchmark Suites for Evaluating Type Inference
+We utilize two benchmark suites StatType-SO and ThaliaType for evaluating LLMs’ type inference performance on code snippets.
 3.1. StatType-SO
-The StatType-SO dataset is prevalent for testing type inference tools (Phan et al., 2018; Saifullah et al., 2020; Dong et al., 2022; Huang et al., 2023b).
+The StatType-SO benchmark suite is prevalent for testing type inference tools (Phan et al., 2018; Saifullah et al., 2020; Dong et al., 2022; Huang et al., 2023b).
 It is constructed from real-world, manually repaired Java code snippets from Stack Overflow.
@@
 The snippets were extracted from 50 randomly selected Stack Overflow posts for each library.
-To prepare the dataset, all parsing errors in the code snippets were first corrected.
+To prepare the benchmark suite, all parsing errors in the code snippets were first corrected.
 Next, the missing import statements were manually inferred and added to the code snippet, serving as the ground truth for evaluation.
-While StatType-SO is a widely adopted dataset for evaluating type inference techniques, it is important to acknowledge the potential for data leakage.
+Figure 4 shows an example from StatType-SO.
+The code snippets from StatType-SO are generally short and make a small number of calls to the libraries and assign the intermediate results to variables.
+While StatType-SO is a widely adopted benchmark suite for evaluating type inference techniques, it is important to acknowledge the potential for data leakage.
 Given its long-standing availability since 2017 at GitHub, StatType-SO may have been inadvertently incorporated into the training data of some LLMs, potentially leading to artificially inflated performance scores.
-3.2. ThaliaType: A New Dataset
-In this paper, we introduce ThaliaType, a new dataset specifically designed to rigorously evaluate the type inference capabilities of LLMs while mitigating the potential biases introduced by data leakage.
-ThaliaType is generated using Thalia, a tool originally developed for testing compilers’ type checkers.
-Recognizing its potential for LLM evaluation, we repurposed Thalia (Sotiropoulos et al., 2024) to generate a diverse set of syntactically valid, well-typed, and type-intensive Java programs.
-Given a set of types, along with the fields and methods defined for each type, Thalia generates programs that utilize a subset of these components.
-Each program consists of a single class with a main method containing variable declarations, field accesses, and method calls.
-Thalia’s ability to generate syntactically correct and well-typed programs, combined with their type-intensive nature, aligns perfectly with the requirements for rigorous type inference evaluation.
-Thalia ensures that the programs are both valid and well-typed, eliminating the need for manual code cleanup or ground truth labeling.
-Furthermore, the type-intensive nature of these programs presents a meaningful challenge for type inference, while the inclusion of field accesses and method calls provides useful context for LLMs during inference.
-To create ThaliaType, we used Thalia to generate 300 code snippets, with the types, fields, and methods drawn from the six libraries used in StatType-SO.
+3.2. ThaliaType: A New Benchmark Suite for Type Inference
+We introduce ThaliaType, a new benchmark suite specifically designed to rigorously evaluate the type inference capabilities of LLMs while mitigating potential biases from data leakage.
+Code snippets in ThaliaType require reasoning over type-related semantic information to accurately infer the types used.
+Through this design, ThaliaType evaluates how effectively LLMs can leverage type information in code snippets, ensuring that the measured performance is generalizable to previously unseen code.
+If an LLM understands the type information provided, then it can correctly infer the types for any given code snippet.
+ThaliaType is generated using Thalia (Chaliasos et al., 2022; Sotiropoulos et al., 2024), a tool originally developed for testing the type system of compilers.
+Thalia has uncovered 117 confirmed or fixed type-related bugs across four major compilers.
+This large number of bugs demonstrates that compiler developers value those found bugs though triggered by automatically generated programs, which indirectly demonstrates the typing relations in Thalia-generated code snippets might resemble the typing relations in real-world code snippets.
+Recognizing its value for LLM evaluation, we repurposed Thalia to create a diverse set of syntactically valid, well-typed Java programs.
+Given a set of types, fields, and methods, Thalia systematically generates programs that utilize a subset of these components.
+By reasoning with the type information within the code snippets, LLMs can infer the types used in ThaliaType.
+The general process for generating ThaliaType code snippets is summarized in Algorithm 1.
+Six libraries that were used in StatType-SO were given to Thalia as input.
+By selecting the same libraries, we ensured that the types in the generated code snippets were familiar to the LLMs, eliminating performance differences caused by varying familiarity with the libraries.
 For each library, 50 code snippets are generated.
-By selecting the same libraries, we ensured that the types in the generated code snippets were familiar to the LLMs, eliminating performance differences based on varying familiarity with the libraries.
-Figure 3 compares the statistics of ThaliaType and StatType-SO.
-The code snippets in both datasets are similar in terms of lines of code and the number of assignments.
+Each snippet is based on a single library, following the convention in StatType-SO.
+In total, 300 code snippets were generated.
+Figure 5 shows an example code snippet generated by ThaliaType.
+Each code snippet consists of a Main class and a test method.
+In this code snippet, the code creates a new LocalTime from a given time represented by a long value of -58.
+The chronology from this LocalTime, which in JodaTime is ISOChronology, is then used to construct the current DateMidnight.
+Finally, the getDayOfYear() method is called to retrieve the ordinal day number of the year.
+Similarity between ThaliaType and StatType-SO.
+The code snippets in ThaliaType are similar to the code snippets from StatType-SO.
+Figure 6 compares the statistics of the two benchmark suites.
+The code snippets in both suites are similar in terms of lines of code and the number of assignments.
 While Thalia-generated programs are type-intensive, ThaliaType includes fewer method calls but utilizes more types (via import statements) in each snippet.
-Our experiments show that SnR achieved comparable performance on both ThaliaType and StatType-SO (section 4.2), demonstrating that ThaliaType provides sufficient semantic information for the type inference task.
-To the best of our knowledge, ThaliaType is the first dataset designed for evaluating LLMs on code snippet type inference while addressing data leakage.
+More importantly, our experiments show that SnR achieved comparable performance on both ThaliaType and StatType-SO (section 5.2), demonstrating that ThaliaType provides as much semantic information as StatType-SO for the type inference task.
+If LLMs genuinely understand the semantics of code snippets, LLMs should be able to perform similarly on both suites.
+To the best of our knowledge, ThaliaType is the first benchmark suite designed for evaluating LLMs on code snippet type inference while addressing data leakage.
 Future research can leverage our replication package to generate additional code snippets using Thalia using their desired libraries to evaluate newer state-of-the-art LLMs.
-4. RQ1: How well do LLMs Perform Type Inference on Unseen Code Snippets
-The data leakage problem is a real threat when evaluating LLM type inference performance, which motivated us to perform a thorough evaluation to uncover their true type inference capabilities.
-To mitigate this issue, in this RQ, we compared LLMs’ type inference performance on the publicly available StatType-SO dataset and the unseen code snippets in ThaliaType, providing insight into the potential impact of data leakage.
+4. Data Leakage in LLM-Based Type Inference
+In this section, we aim to investigate whether the evaluation of LLM-based type inference techniques may be affected by the data leakage issue (Inan et al., 2021).
+Data leakage occurs when the training datasets overlap with the benchmark datasets used for evaluation, leading models to achieve artificially high performance by memorizing examples rather than demonstrating genuine generalization.
+The evaluation of the prior studies that either compared against (Huang et al., 2023a; Chen et al., 2024) or incorporated LLMs (Kabir et al., 2025) largely relied on the StatType-SO dataset as the benchmark suite, which has been publicly available on GitHub since 2017 (Sta, 2017).
+Because the GPT-based models used in these evaluations were likely trained on datasets containing StatType-SO, their reported performance may be affected by data leakage.
+However, directly detecting data leakage from LLMs’ training is challenging, as most datasets are proprietary and undisclosed, with only a few released to the broader community.
+Two widely used LLM families, GPT (OpenAI, 2024) and Llama (Meta, 2024), are trained on proprietary datasets with only partial disclosure of their training data.
+GPT’s creator, OpenAI, has stated that the models are trained on a mixture of publicly available data (such as internet text) and data licensed from third-party providers.
+However, the exact composition of the dataset remains undisclosed, and access to the model is limited to APIs.
+Similarly, while Llama models are open-weight and thus freely available for download and use, the data used to train them is not publicly released.
+Meta discloses that Llama is trained on publicly available data, primarily sourced from the internet, but the exact sources are not specified.
+This lack of transparency makes it difficult to assess the risk of data leakage in either model.
+In contrast, StarCoder2 (Lozhkov et al., 2024) is trained entirely on the openly available dataset StackV2 (BigCode, 2024), which comprises 67.5TB of source code across 658 programming languages.
+Because this dataset is openly accessible, it is possible to directly examine the training corpus and assess the extent of any data leakage.
+| Repository Owner | Repository Name | # of code snippets | Last Commit Date |
+| miketran238 | AndroidOracle | 50 | 2017-02-14 |
+| miketran238 | JodatimeOracle | 50 | 2017-02-12 |
+| mrthlinh | Oracle-GWT | 50 | 2017-02-13 |
+| mrthlinh | Oracle-Hibernate | 50 | 2017-02-13 |
+| mrthlinh | TypeResolution_Oracle | 219 | 2017-02-12 |
+| pdhung3012 | TypeResolution_Oracle | 267 | 2017-06-30 |
+To investigate the content of the StackV2, the metadata of the files are examined.
+Using the matching procedure described in Algorithm 2, file names in StackV2 were compared against those in the StatType-SO benchmark suite, which identified repositories containing overlapping code snippets.
+This process produced 26 repositories and three repository owners with five or more matching file names, forming the basis for manual inspection.
+Manual analysis identified six repositories (shown in Table 2) containing code snippets alongside their ground truth from the StatType-SO benchmark suite.
+Notably, the TypeResolution_Oracle repository by mrthlinh includes 219 code snippets with ground truth, excluding only those dependent on the Android library.
+In addition, pdhung3012’s TypeResolution_Oracle repository contains the complete StatType-SO benchmark suite.
+All identified repositories were committed in 2017, well before the knowledge cutoffs of GPT and Llama, making it highly likely that these models used for evaluation were trained on the full StatType-SO benchmark suite.
+5. RQ1: How well do LLMs perform type inference on unseen code snippets?
+Given the strong possibility of data leakage, we sought to rigorously assess the genuine type inference capabilities of LLMs.
+To do this, RQ1 investigates their performance on two benchmark suites: the original StatType-SO (which may have been seen during training), and ThaliaType (which contains only unseen code snippets).
+This comparison allows us to isolate the impact of potential training overlap and better understand how well these models generalize to genuinely novel code.
 5.1. Method
-Code snippets from StatType-SO and ThaliaType are given to LLMs using the prompt pattern in Figure 1 and also our baseline SnR one at a time without import statements.
+Code snippets from StatType-SO and ThaliaType are given to LLMs using the prompt pattern in Figure 3 and also our baseline SnR, one at a time, without import statements.
 To check the output, import statements in LLMs’ response are extracted and compared against the original import statements in the code snippets.
 To ensure reproducibility, all the inferences are performed with a fixed seed (set to one) and a temperature of zero.
-GPT-4o-mini and GPT-4o are accessed using the OpenAI API (OPE, 2024b), while Llama3.1:8b and Llama3.1:70b are accessed through a self-hosted Ollama API (OLL, 2024) running on an A100 GPU.
-The precision, recall, and F1-scores are computed to evaluate the performance of type inference.
+GPT-4o-mini and GPT-4o are accessed via the OpenAI API (OPE, 2024b), while StarCoder2:15b, Llama3.1:8b, and Llama3.1:70b are accessed through a self-hosted Ollama API (OLL, 2024) running on an A100 GPU.
+The total cost of using the GPT models for RQ1 was $4.47, based on OpenAI’s pricing as of early 2025.
+Precision, recall, and F1-scores are computed to evaluate the performance of type inference.
 Precision measures the proportion of correctly inferred FQNs among all inferred FQNs, while recall measures the proportion of expected FQNs that are correctly inferred.
 5.2. Results
-LLMs achieved lower performance on unseen code snippets in ThaliaType despite their strong results in StatType-SO (Figure 4).
-Specifically, as shown in Figure 4(b), on the unseen ThaliaType code snippets, GPT-4o achieved a precision and recall of 54.74% and 44.54% respectively, a 42.8% and 53.1% decrease in precision and recall compared to results on StatType-SO.
-SnR outperformed all LLMs in ThaliaType with a precision of 84.15% and a recall of 84.43% .
-Even on StatType-SO, SnR achieved second best precision and recall of 95.50% and 91.46% .
-The large disparity between the LLMs’ performance on StatType-SO and unseen code snippets in ThaliaType suggests that LLMs are highly likely affected by data leakage which potentially inflated LLMs’ type inference performance on StatType-SO.
+Despite achieving strong results in StatType-SO (Figure 7), all evaluated LLMs demonstrated lower performance on unseen code snippets in ThaliaType.
+StarCoder2:15b consistently ranked between Llama3.1:8b and Llama3.1:70b across both benchmark suites.
+Specifically, as shown in Figure 7(a), StarCoder2:15b achieved an F1 score of 81.67% on StatType-SO, higher than Llama3.1:8b, but lower than Llama3.1:70b, GPT-4o, and GPT-4o-mini.
+However, its performance dropped sharply to 27.08% on ThaliaType, representing a 66.8% decrease.
+Despite this decline, it still outperformed Llama3.1:8b but remained below the other larger models.
+Notably, even the best performing model, GPT-4o, experienced a 48.5% decrease in F1 score when evaluated on ThaliaType.
+The consistent performance drop across all models, including StarCoder2:15b with confirmed data leakage, suggests that LLMs are likely affected by data leakage, which potentially inflated LLMs’ type inference performance on StatType-SO.
 Unlike LLMs, SnR relies solely on the names of types, method names, and relationships between types in code snippets.
-The fact that LLMs outperformed SnR on StatType-SO suggests that LLMs leverage additional information not used by SnR to perform type inference.
-Conversely, the poor performance of LLMs on ThaliaType compared to SnR implies that LLMs failed to fully exploit the simple names, method calls, and type relations within the code snippet.
+SnR only experienced a 9.8% decrease in F1 score when evaluated on ThaliaType, due to the difference in available information in the code snippets.
+The fact that LLMs outperformed SnR on StatType-SO while performing poorly on ThaliaType suggests that LLMs leverage additional information in StatType-SO to perform type inference.
 These performance discrepancies motivated further investigations into the specific factors contributing to LLMs’ type inference capabilities.
-In particular, we explore in RQ2 (section 5) if LLMs leverage the available semantic information in code snippets or simply rely on matching exact syntax of the code snippets from training data.
+In particular, in RQ2 (section 6), we investigate what types of information LLMs might leverage for type inference, and whether their performance depends on matching exact syntax of the code snippets seen during training.
 | Document Frequency | [0,1e2) | [1e2,1e3) | [1e3,1e4) | [1e4,1e5) | ¿=1e5 | ||||||
@@
 | SnR | 27 | 100.00% | 60 | 86.96% | 297 | 95.19% | 372 | 86.71% | 433 | 93.52% | |
+| StarCoder2:15b | 7 | 25.93% | 36 | 52.17% | 235 | 75.32% | 372 | 86.71% | 404 | 87.26% | |
 | Llama3.1:8b | 2 | 7.41% | 28 | 40.58% | 173 | 55.45% | 322 | 75.06% | 378 | 81.64% | |
@@
 | SnR | 735 | 73.94% | 533 | 89.28% | 407 | 92.71% | 300 | 92.59% | 292 | 88.22% | |
+| StarCoder2:15b | 22 | 2.21% | 58 | 9.72% | 94 | 21.41% | 120 | 37.04% | 234 | 70.69% | |
 | Llama3.1:8b | 19 | 1.91% | 42 | 7.04% | 64 | 14.58% | 138 | 42.59% | 258 | 77.95% | |
@@
 5.2.1. Impact of Document Frequency on LLM Type Inference
-To better understand why there is a large disparity between StatType-SO and ThaliaType, we investigated how the document frequency of a type impacts the performance of LLMs.
+To better understand the large performance disparity between StatType-SO and ThaliaType, we investigated how the document frequency of a type impacts the performance of LLMs.
 Document frequency measures the number of source files that reference a specific type.
-Since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024), it is plausible that their performance may vary depending on the prevalence of a type within a LLMs’ training data.
+Since LLM outputs are biased by training data (Huang et al., 2024; Liu et al., 2024), their performance may vary depending on the prevalence of a type in the training corpus.
 Understanding this bias helps ensure that LLMs deliver similar performance across all potential types used in the real world.
@@
 Specifically, the document frequency is calculated as follows for a type .
-Based on the document frequency, we categorized the FQNs in the StatType-SO and ThaliaType datasets into distinct frequency ranges.
+Based on the document frequency, we categorized the FQNs in the StatType-SO and ThaliaType benchmark suites into distinct frequency ranges.
 For each group, we calculated the recall to investigate how a type’s real-world frequency impacts LLM performance.
@@
 Results.
-LLMs performed poorly on less frequently used FQNs and worse on the unseen dataset ThaliaType than on StatType-SO, even for frequently used FQNs.
-To analyze how performance varies across FQNs with differing levels of usage frequencies, we grouped FQNs into frequency ranges by orders of magnitude (0-100, 100-1,000, 1,000-10,000, 10,000-100,000, and over 100,000 GitHub files).
-For the least frequently used FQNs in ThaliaType, GPT-4o only achieved a recall of 22.22% on StatType-SO and 10.36% on ThaliaType.
+LLMs performed poorly on less frequently used FQNs and worse on the unseen benchmark ThaliaType than on StatType-SO, even for frequently used FQNs.
+To analyze how performance varies across FQNs with differing usage frequencies, we grouped FQNs into frequency ranges by orders of magnitude (0-100, 100-1,000, 1,000-10,000, 10,000-100,000, and over 100,000 GitHub files).
+The results, shown in Table 3, indicate that LLM performance degrades as the document frequency of FQNs decreases.
+Specifically, for the least frequently used FQNs, GPT-4o only achieved a recall of 22.22% on StatType-SO and 10.36% on ThaliaType.
 In contrast, SnR demonstrated consistent performance regardless of a type’s popularity, achieving recalls of 100.00% and 73.94% on StatType-SO and ThaliaType respectively.
 On ThaliaType, LLMs only outperformed SnR on the most frequently used FQNs, specifically those appearing in over 100,000 GitHub source files.
-This disparity may arise from LLMs preferring more frequent FQN for a given simple name regardless of the context in the code snippet.
+This disparity may arise from LLMs preferring more frequent FQNs for a given simple name regardless of the context in the code snippet.
 For example, for a type with the simple name View, if android.view.View is the most frequent type in the ¿=1e5 category and javax.swing.text.View belongs to a less frequent category, then always recommending android.view.View will result in a high recall in the ¿=1e5 category but a lower recall in other categories.
-5. RQ2: To what extent do LLMs understand the semantics of code snippets?
-The performance decline on unseen data observed in RQ1 highlights the need to investigate how LLMs handle type inference.
+Interestingly, despite StarCoder2:15b being trained on code snippets from StatType-SO, it achieved a maximum recall of only 87.26% .
+As in the previous analysis, its performance generally fell between Llama3.1:8b and Llama3.1:70b on both StatType-SO and ThaliaType.
+Although training data may inflate performance, StarCoder2:15b does not reproduce code snippets verbatim, as models are designed to generalize beyond specific examples.
+This generalization, while desirable, makes detecting instances of data leakage difficult.
+6. RQ2: To what extent do LLMs understand the execution semantics of code snippets?
+The performance decline on unseen data observed in RQ1 highlights the need to investigate how LLMs perform type inference.
 To explore whether this decline stems from a lack of semantic understanding in LLMs, we conducted an experiment using semantic-preserving transformations that alter syntax without changing the semantics.
-Our hypothesis is that if LLMs possess a deep understanding of the code snippets’ semantics, their performance should remain stable despite such transformations.
+Our hypothesis is that if LLMs possess a deep understanding of code snippet semantics, their performance should remain stable despite such transformations.
 Conversely, significant performance degradation would suggest that LLMs rely heavily on superficial syntactic patterns rather than a genuine semantic comprehension.
@@
 The Wilcoxon signed-rank test (Wilcoxon, 1992) was used to determine whether the transformations caused significant changes in precision, recall, and F1 scores.
+The average running cost per transformation for GPT models was $6.30 in RQ2, slightly higher than in RQ1 because of longer code snippets.
 6.1.1. Background: Java Grammar
@@
 Classes may contain fields, methods, and blocks, with methods comprising blocks of statements and expressions.
-For simplicity, modifiers (e.g., public, static) were excluded, as they remain unchanged in the transformations.
-Other Java statements and expressions were similarly omitted as they follow the same transformation principles.
+Highly repetitive Java statements and expressions were omitted as they follow the same transformation principles.
 6.1.2. Identifier Renaming
 This transformation investigates whether LLMs rely on specific identifier names for type inference.
+Although identifier names can provide useful signals, overreliance on unique identifiers or specific identifier sequences seen during training may limit a model’s ability to generalize.
+In real-world code snippets, identifiers are often lowercase variants of type names or abbreviated forms, providing no additional semantic information.
 The RenameIdentifier function in Algorithm 3 details the process, which systematically renames identifiers such as variables, methods, classes, and packages in the code snippet.
@@
 Next, the algorithm processes method declarations.
-To avoid renaming references that could alter semantics, all the method call expressions (s with the form ) that have the first optional part () is potentially external, thus added to a skip_list (Algorithm 1-1) to be filtered out (Algorithm 1).
+To avoid renaming references that could alter semantics, all the method call expressions (s with the form ) that have the first optional part () are potentially external, thus added to a skip_list (lines 3-3) to be filtered out (Algorithm 3).
 Additionally, overridden methods annotated with @Override are skipped, as their names must match those in the super class (Algorithm 3).
-Lastly, the algorithm renames classes and packages (lines 1-1), following the same procedure used for variables.
+Lastly, the algorithm renames classes and packages (lines 3-3), following the same procedure used for variables.
 The details are omitted here for brevity.
@@
 Figure 9 provides an example of this transformation, showing how variables, methods, classes, and packages are renamed.
-Method m which likely overrides the method m on type a is preserved out of an abundance of caution to preserve semantic consistency.
+The method m declared at line 5 likely overrides a method on the super class, A.m().
+Out of an abundance of caution to preserve semantic consistency, m is not renamed.
 Furthermore, it should be noted that different types of names are renamed separately.
-For example, the method name n is renamed to a different name than the variable name n.
+For example, the method name n is renamed differently from the variable name n.
 6.1.3. Code Lowering
 The code lowering transformation examines the impact of structural changes on LLMs’ type inference.
+A key aspect of type inference is understanding the types present in a program and how they relate to one another.
+Code lowering does not alter these type relationships but instead disrupts the exact sequence of tokens presented to the model.
+If a model relies heavily on specific token sequences seen during training, its generalizability may be impaired, leading to failures in type inference when the same program is presented in a different structure.
 The algorithm first extracts and traverses all the expressions (Algorithm 3) in the code snippet.
-For each expression, if its parent is not an expression statement (s with the form ) and it is not a loop condition (lines 1-1), the algorithm creates a new variable with a random name, assigns the expression to the variable, and inserts this assignment statement before the expression (lines 1-1).
+For each expression, if its parent is not an expression statement (s with the form ) and it is not a loop condition (lines 3-3), the algorithm creates a new variable with a random name, assigns the expression to the variable, and inserts this assignment statement before the expression (lines 3-3).
 Next, the original expression is replaced with the newly created variable assigned with the value of the original expression (Algorithm 3).
@@
 In addition to expressions, code lowering also applies to fields with an initializer.
-For each field in the code snippet (Algorithm 1), if it has an initializer (Algorithm 1), the algorithm splits it into a declaration and a block initializing the field (lines 1-1).
+For each field in the code snippet (Algorithm 3), if it has an initializer (Algorithm 3), the algorithm splits it into a declaration and a block initializing the field (lines 3-3).
 Figure 10 demonstrates this transformation.
 In this example, the field C c = null is split into a declaration and an initializer block with c = null.
-Additionally, the expression in the if statement is replaced with a newly created variable, and an assignment statement is inserted before the expression’s usage to assign its original value to the variable.
+Additionally, the expression in the if statement is replaced with a newly created variable, and an assignment statement is inserted before its use to assign the original value to that variable.
 6.1.4. Adding Keyword Comments
 The keyword comment transformation evaluates whether comments containing Java keywords distract LLMs.
-AddKeyword function in Algorithm 1 appends a single Java keyword (Gosling et al., 2023) as a line comment at the end of each line in the code snippet.
+If a model relies on specific sequences of keywords seen during training, rather than on the actual content of the code, then this transformation may lead to reduced performance and an inability to generalize to unseen code snippets.
+To perform this transformation, the AddKeyword function in Algorithm 3 appends a single Java keyword (Gosling et al., 2023) as a line comment at the end of each line in the code snippet.
 An example of the transformed snippet is presented in Figure 11.
-These comments do not alter the semantics of the code but provide additional syntactic noise.
+6.1.5. Putting Everything Together
+Figure 12 illustrates the complete set of transformations used in RQ2.
+Each of the first three transformations was applied independently to measure its individual impact.
+The final transformation combines all three sequentially to more strongly perturb the input, thereby making it more difficult for models to recall StatType-SO code snippets, if they were seen during training.
+These transformed code snippets are then passed to LLMs.
+The inferred types are analyzed to see if transformations significantly impacted LLMs understanding of the types used in the code snippet, affecting performance.
 6.2. Results
-| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
-| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
-| StatType-SO | 95.50% | 91.46% | 93.44% | 76.92% | 69.46% | 73.00% | 86.08% | 83.69% | 84.87% | 86.34% | 89.92% | 88.09% | 95.66% | 95.00% | 95.33% |
-| Id Renaming | 95.50% | 91.46% | 93.44% | 65.28% | 56.69% | 60.68% | 85.34% | 76.15% | 80.49% | 89.63% | 88.46% | 89.04% | 97.03% | 95.62% | 96.32% |
-| Code Lowering | 95.43% | 91.62% | 93.49% | 69.43% | 65.15% | 67.22% | 83.81% | 80.46% | 82.10% | 86.14% | 88.92% | 87.51% | 94.82% | 95.69% | 95.25% |
-| Comment Adding | 95.50% | 91.46% | 93.44% | 78.65% | 67.46% | 72.63% | 85.39% | 80.92% | 83.10% | 87.90% | 89.38% | 88.63% | 96.06% | 95.77% | 95.92% |
-| All | 95.43% | 91.62% | 93.49% | 55.83% | 47.15% | 51.13% | 77.93% | 70.08% | 73.80% | 84.60% | 82.85% | 83.72% | 93.29% | 94.15% | 93.72% |
-| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
-| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
-| ThaliaType | 84.15% | 84.43% | 84.29% | 31.27% | 19.40% | 23.95% | 61.58% | 25.85% | 36.41% | 66.64% | 37.73% | 48.18% | 54.74% | 44.54% | 49.12% |
-| Id Renaming | 84.15% | 84.43% | 84.29% | 31.36% | 19.11% | 23.74% | 57.08% | 24.47% | 34.25% | 52.47% | 38.44% | 44.37% | 43.88% | 45.77% | 44.80% |
-| Code Lowering | 84.14% | 84.39% | 84.27% | 27.06% | 18.21% | 21.77% | 60.16% | 25.25% | 35.57% | 58.04% | 36.01% | 44.45% | 43.45% | 44.43% | 43.93% |
-| Comment Adding | 84.15% | 84.43% | 84.29% | 25.00% | 19.22% | 21.73% | 59.56% | 25.07% | 35.28% | 64.27% | 36.57% | 46.62% | 32.26% | 44.80% | 37.51% |
-| All | 84.14% | 84.39% | 84.27% | 27.57% | 17.77% | 21.61% | 45.92% | 23.87% | 31.41% | 47.50% | 36.87% | 41.52% | 56.30% | 45.29% | 50.20% |
-The effects of applying transformations to the StatType-SO and ThaliaType datasets are shown in Tables 3(a) and 3(b), respectively.
-These transformations significantly impacted the performance of all LLMs, resulting in significant decreases in both precision and recall.
-When all three transformations were applied simultaneously, the decrease in performance for LLMs was amplified.
-The F1-score dropped significantly () for all LLMs on StatType-SO (e.g., 1.7% decrease on GPT-4o, and 30.0% decrease on Llama3.1:8b).
-Precision dropped significantly for all LLMs across both the StatType-SO and ThaliaType datasets, while recall also declined on both datasets for nearly all models.
-Comparing the F1-score changes revealed that LLMs were more affected by the transformations in StatType-SO than in ThaliaType.
-Unlike ThaliaType, which consists entirely of unseen code snippets, the more substantial decline in StatType-SO indicates potential data leakage, where LLMs rely on memorized patterns rather than demonstrating a deep semantic understanding of the code.
+The effects of applying transformations to the StatType-SO and ThaliaType benchmark suites are shown in sections 6.2 and 6.2, respectively.
+These transformations significantly affected the precision, recall, and F1-scores of all evaluated LLMs.
+However, the results are complex and merit detailed examination.
+| SnR | StarCoder2:15b | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||||
+| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
+| StatType-SO | 95.50% | 91.46% | 93.44% | 82.28% | 81.08% | 81.67% | 76.92% | 69.46% | 73.00% | 86.08% | 83.69% | 84.87% | 86.34% | 89.92% | 88.09% | 95.66% | 95.00% | 95.33% |
+| Id Renaming | 95.50% | 91.46% | 93.44% | 82.09% | 71.23% | 76.28% | 65.28% | 56.69% | 60.68% | 85.34% | 76.15% | 80.49% | 89.63% | 88.46% | 89.04% | 97.03% | 95.62% | 96.32% |
+| Code Lowering | 95.43% | 91.62% | 93.49% | 83.72% | 79.92% | 81.78% | 69.43% | 65.15% | 67.22% | 83.81% | 80.46% | 82.10% | 86.14% | 88.92% | 87.51% | 94.82% | 95.69% | 95.25% |
+| Comment Adding | 95.50% | 91.46% | 93.44% | 83.98% | 75.00% | 79.24% | 78.65% | 67.46% | 72.63% | 85.39% | 80.92% | 83.10% | 87.90% | 89.38% | 88.63% | 96.06% | 95.77% | 95.92% |
+| [gray]0.8 black All | 95.43% | 91.62% | 93.49% | 79.34% | 59.38% | 67.93% | 55.83% | 47.15% | 51.13% | 77.93% | 70.08% | 73.80% | 84.60% | 82.85% | 83.72% | 93.29% | 94.15% | 93.72% |
+| SnR | StarCoder2:15b | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||||
+| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
+| ThaliaType | 84.15% | 84.43% | 84.29% | 43.46% | 19.66% | 27.08% | 31.27% | 19.40% | 23.95% | 61.58% | 25.85% | 36.41% | 66.64% | 37.73% | 48.18% | 54.74% | 44.54% | 49.12% |
+| Id Renaming | 84.15% | 84.43% | 84.29% | 46.24% | 22.42% | 30.20% | 31.36% | 19.11% | 23.74% | 57.08% | 24.47% | 34.25% | 52.47% | 38.44% | 44.37% | 43.88% | 45.77% | 44.80% |
+| Code Lowering | 84.14% | 84.39% | 84.27% | 48.09% | 23.87% | 31.91% | 27.06% | 18.21% | 21.77% | 60.16% | 25.25% | 35.57% | 58.04% | 36.01% | 44.45% | 43.45% | 44.43% | 43.93% |
+| Comment Adding | 84.15% | 84.43% | 84.29% | 42.96% | 19.66% | 26.98% | 25.00% | 19.22% | 21.73% | 59.56% | 25.07% | 35.28% | 64.27% | 36.57% | 46.62% | 32.26% | 44.80% | 37.51% |
+| [gray]0.8 black All | 84.14% | 84.39% | 84.27% | 40.81% | 20.74% | 27.51% | 27.57% | 17.77% | 21.61% | 45.92% | 23.87% | 31.41% | 47.50% | 36.87% | 41.52% | 56.30% | 45.29% | 50.20% |
+Focusing on each transformation in isolation (identifier renaming, code lowering, comment adding), the results show that LLMs generally exhibit resilience, meaning that simple transformations often do not significantly change a model’s performance.
+This was particularly true for more advanced models such as GPT-4o but also held for smaller models like StarCoder2:15b.
+Notably, StarCoder2:15b, known to have been trained on StatType-SO code snippets, exhibited significant performance drops only under identifier renaming (recall) and comment adding (recall and F1-score) when evaluated on StatType-SO.
+The isolated transformation results on StatType-SO and ThaliaType indicate a degree of generalizability in LLMs’ understanding of code.
+However, when all three transformations were applied simultaneously, performance changes were amplified, clearly, significantly, and negatively affecting all LLMs’ precision, recall, and F1-scores on StatType-SO.
+For example, F1-scores dropped by 1.7% on GPT-4o, 16.8% on StarCoder2:15b, and 30.0% on Llama3.1:8b.
+Similar trends were observed for precision and recall.
+In contrast, applying the same three transformations on ThaliaType code snippets produced less consistent and generally smaller effects.
+In some cases, performance even improved.
+For instance, StarCoder2:15b achieved a higher F1-score despite a lowered precision, and GPT-4o demonstrated higher precision, with no significant changes in other metrics on ThaliaType code snippets.
+Focusing on GPT-4o on StatType-SO, despite identifier renaming and comment adding causing insignificant performance increases, and code lowering causing a small but significant decrease in precision without impacting F1, when all combined, significantly decreased performance in precision, recall, and F1-scores.
+These effects were not observed for any model using ThaliaType code snippets.
+The discrepancies in performance indicate that the results in StatType-SO may be influenced by data leakage and have limited generalizability compared to the results in ThaliaType.
+A potential explanation is that transformations reduced the ability of LLMs to recognize code patterns from StatType-SO snippets from training, leading to degraded performance.
+In contrast, ThaliaType consists of newly generated code snippets that were not in the training data.
+As a result, models must rely more on reasoning over execution semantics rather than recall from training.
+The finding that performance on ThaliaType does not consistently decline, and sometimes even improves, suggests that current LLMs can withstand simple, execution-semantic preserving transformations.
+LLMs generalize more effectively on ThaliaType than their StatType-SO performance suggests.
 In contrast, SnR, which analyzes the semantic meaning behind the code snippets, was unaffected by these transformations, experiencing only minor, statistically insignificant variations under the code lowering transformation.
 These variations occurred when multiple types satisfied all constraints, with the order of type variables serving as a tie-breaker.
-6. RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
-Building upon the findings of RQ2, which demonstrated the sensitivity of LLMs to subtle syntactic variations, we investigate the fundamental sources of LLM’ type inference performance.
-Specifically, we aim to identify the minimal set of syntax elements sufficient for LLMs to perform type inference.
-6.1. Method
-We applied delta debugging to systematically minimize the input code snippet while preserving consistent inference results.
-This approach enables us to isolate the exact syntactic elements that are crucial for LLMs’ type inference.
-The intuition is that tokens conveying more semantic information and critical for the type inference task, such as simple names, should be preserved, whereas tokens carrying less semantic information, such as whitespaces, can be removed.
-Identifying these crucial syntactic elements provides valuable insights into the factors that influence LLM performance in type inference and can inform the development of more robust and efficient type inference systems.
-6.1.1. Background: Delta Debugging
-Delta debugging (Zeller and Hildebrandt, 2002), specifically the ddmin algorithm, is a widely used technique in software debugging to minimize a given input while preserving a certain property.
-In our study, the property we intend to preserve is that the minimized code snippet contains sufficient information for LLMs to infer the same import statements as the original code snippet.
-The minimized code snippet preserves this property, representing the elements in the code snippet sufficient for LLMs to conduct type inference.
-In this research question, in addition to ddmin, we also used Perses (Sun et al., 2018) to minimize code snippets due to its high efficiency.
-Notably, Perses achieves this efficiency by ensuring the transformed program remains syntactically valid.
-However, since LLMs can interpret syntactically invalid programs, we apply ddmin to further minimize the output from Perses.
-Unlike in previous RQs, the purpose of the minimization here is to identify the exact syntax elements required for LLMs’ type inference, meaning that syntactical validity does not need to be preserved.
-In the end, the minimized program is 1-minimal—meaning that the removal of any single token will cause the type inference tool to fail in inferring the same import statements.
-6.1.2. Code Snippet Minimization
-Algorithm 2 outlines the core conceptual workflow of code snippet minimization.
-The minimization process follows a trial-and-error approach as defined by the Minimize function.
-During this process, a minimal snippet min is maintained and initially set to the original snippet (Algorithm 2).
-Given the current minimal snippet min, a delta debugging algorithm generates a set of smaller snippets that may retain the property, i.e., the LLM infers the same import statements as the original ones for the snippet (Algorithm 2).
-The primary differences among different delta debugging algorithms lay in the strategy of generating smaller snippets.
-For brevity, we abstract the distinctions between the delta debuggers, (i.e., ddmin and Perses) using the ComputeSmallerSnippets function.
-If any of the smaller snippets preserve the property, the current minimal snippet min is updated, and the process continues (lines 2-2).
-Otherwise, if none of the smaller snippets preserve the property, indicating that no further progress can be made, the minimization terminates (lines 2-2).
-In our experiment, to minimize a code snippet, the import statement inferred by the LLM is first extracted (Algorithm 2), after which, the algorithm applies snippet minimization with both Perses and ddmin to derive the final minimal snippet (lines 2-2).
-Figure 9 illustrates this process using a code snippet from StatType-SO (Figure 9(a)).
-In this example, GPT-4o-mini initially inferred four types: Session, Transaction, SessionFactory, and Configuration (Figure 9(b)).
-After applying Algorithm 2, the snippet was minimized to a single line ibernate (Figure 9(c)).
-The minimization process is computationally expensive, particularly the repeated type inference performed for each smaller variant (Algorithm 2).
-The runtime for each type inference operation varies depending on the model’s size and complexity.
-With tens of thousands of inference calls required, due to cost reasons, we focused on GPT-4o-mini only for this research question.
-6.2. Result
-The minimization process eliminated a greater portion of tokens from code snippets in StatType-SO compared to ThaliaType, as summarized in Figure 10.
-On average, only 6.37% of the tokens were retained for each code snippet in StatType-SO, whereas 14.73% of the tokens were preserved for each code snippet in ThaliaType, more than double.
-While the original datasets contained 95803 and 91166 tokens for StatType-SO and ThaliaType respectively, the minimized versions only contained 4619 and 16317 tokens (Table 4).
-The higher removal ratio indicates that many of the tokens in the original code snippet are not required to achieve the high type inference performance seen in StatType-SO.
-| StatType-SO | ThaliaType | |||||
-| Category | Ratio | # in Original | # in Minimized | Ratio | # in Original | # in Minimized |
-| Average | 4.82% | 95803 | 4619 | 17.90% | 91166 | 16317 |
-| Control and Structural | 3.17% | 36791 | 1165 | 14.39% | 37864 | 5450 |
-| Inferred Simple Name | 16.88% | 4954 | 836 | 53.33% | 4478 | 2388 |
-| Whitespace | 6.18% | 29388 | 1816 | 20.95% | 29144 | 6105 |
-| Other | 3.25% | 24670 | 802 | 12.06% | 19680 | 2374 |
-To analyze how different types of tokens are affected by the minimization process, tokens are categorized into four groups, control and structural, whitespace, inferred simple name, and other.
-The control and structural category includes symbols (e.g., parenthesis, semicolon), operators (e.g.
-plus, minus, equals), and keywords (e.g.
-if, while), as defined by the Java Language Specification (JLS, 2021).
-The whitespace category consists of tokens composed entirely of whitespace characters (e.g., space, new line, tab).
-The inferred simple name category includes simple names of the types inferred by the LLM.
-If the simple name is not inferred, it falls into the other category, which also includes tokens such as integer or floating-point literals.
-As shown in Table 4 control and structural tokens, along with those in the other category, are frequently removed during minimization.
-In contrast, whitespace and inferred simple name tokens are often retained.
-We further analyzed the inferred simple name category to determine whether simple names were consistently preserved in minimized code snippets.
-Retaining these names is crucial for the task of inferring FQNs of types, since such types are often referenced by their simple names.
-However, as shown in Figure 11, these simple names are frequently absent in minimized versions.
-To quantify this effect, we calculated the percentage decrease in the unique simple names after minimization.
-To account for simple names that were not inferred, we only included those for which LLMs provided an FQN.
-For example, while the code snippet in Figure 9(a) contained all four of the inferred type’s simple names (Figure 9(b)), none were preserved in the minimized version (Figure 9(c)), representing a 100% decrease.
-On average, the number of unique simple names in StatType-SO decreased by 41.49% , suggesting that inferring the types in StatType-SO often does not require simple names.
-In contrast, ThaliaType exhibited only a 9.12% decrease.
-These findings suggest that LLMs may rely more on direct recollection of their training data than truly interpreting the semantic information in code snippets in StatType-SO.
-Since type inference inherently requires the simple names to resolve FQNs, the ability of LLMs to infer types without these simple names indicates that they may not be solving the task as intended.
-Instead, their high performance may stem from exploiting superficial patterns in code snippets.
-This highlights a weakness in the StatType-SO dataset: its potential predictability and insufficient penalization of overfitting.
-Consequently, LLM performance on StatType-SO may not generalize effectively to unseen code snippets.
-To ensure a more robust and generalizable assessment of LLM performance, future work should focus on evaluating models using datasets composed entirely of unseen code snippets.
 7. Threats to Validity
-There are three main threats to validity.
-First, variations in the document frequency of FQNs between ThaliaType and StatType-SO datasets could influence LLM type inference performance, since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024).
+There are four main threats to validity.
+First, the representativeness of ThaliaType may not fully capture certain qualities of real-world Java code snippets, such as variable naming.
+However, this limitation is mitigated by the primary objective of our evaluation, which is to assess the type inference capabilities of LLMs.
+Specifically, ThaliaType is designed to evaluate whether LLMs can infer correct type information by reasoning about the execution semantics of the code snippets, rather than merely recalling examples from training.
+We acknowledge that LLMs may leverage additional semantic cues present in StatType-SO but not found in ThaliaType for type inference.
+To account for these potential differences, RQ2 applies transformations to both benchmark suites.
+The identifier renaming transformation on StatType-SO even slightly increased type inference performance of GPT-4o, which means that factors such as identifier names do not significantly affect GPT-4o’s performance on StatType-SO and cannot explain LLMs’ diminished performance on ThaliaType.
+Furthermore, the strong performance of SnR on both StatType-SO and ThaliaType indicates that, (1) ThaliaType shares key features that are necessary for type inference with StatType-SO, (2) it presents a meaningful challenge for type inference techniques, and (3) it exposes areas where LLMs require further improvement.
+Second, variations in the document frequency of FQNs between ThaliaType and StatType-SO could influence LLM type inference performance, since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024).
 LLM’s performance may vary depending on the prevalence of a type within LLM’s training data.
-To address this, we analyzed LLM performance across different document frequency levels.
+To address this, we analyzed the performance of LLMs across different document frequency levels.
 Despite these efforts, LLMs consistently exhibited lower performance on unseen code snippets from ThaliaType compared to StatType-SO.
-Second, our findings may not generalize to other LLMs.
-We mitigated this by following best practices (Sallou et al., 2024) and evaluated a diverse set of models, including both open weight (Llama3.1:8b, Llama3.1:70b) and the state-of-the-art closed models (GPT-4o, GPT-4o-mini) of varying sizes, providing a representative sample of current LLM capabilities.
-However, due to the high cost of performing minimization, we were only able to evaluate GPT-4o-mini for RQ3.
-Nevertheless, since all current LLMs are based on transformers, their underlying type inference mechanisms operate in a similar manner.
-Thus our findings in RQ3 remain generalizable—specifically, that the StatType-SO dataset is insufficient for evaluating type inference.
-Third, prompt engineering may enable LLM to achieve better performance for type inference on code snippets.
-To mitigate this threat, we followed best practices to design a simple but effective prompt (LLM, 2024) which achieved high performance on the StatType-SO dataset.
+Third, our findings may not generalize to other LLMs.
+We mitigated this by following best practices (Sallou et al., 2024) and evaluated a diverse set of models, including both open-weight (StarCoder2:15b, Llama3.1:8b, Llama3.1:70b) and state-of-the-art closed models (GPT-4o, GPT-4o-mini) of varying sizes, providing a representative sample of current LLM capabilities.
+Fourth, prompt engineering may enable LLM to achieve better performance for type inference on code snippets.
+To mitigate this threat, we followed best practices to design a simple but effective prompt (LLM, 2024), which achieved high performance on the StatType-SO suite.
 Regardless, data leakage remains a fundamental issue that can influence LLM evaluations regardless of prompt quality.
@@
 8. Related Work
-In this section, we discuss related work in the fields of type inference for code snippets and LLM data leakage.
+In this section, we briefly discuss different techniques for type inference on code snippets and LLM data leakage in other fields including automated program repair, code generation, refactoring, and more.
 8.1. Type Inference for Code Snippets
@@
 Machine-Learning-Based Approaches.
-Machine-learning-based approaches typically leverage a model trained with a large set of programs from open-sourced projects.
+Machine-learning-based approaches typically leverage models trained on a large set of programs from open-source projects.
 Phan et al.
@@
 With such knowledge, StatType can derive the FQN for an API based on the neighboring API names.
-A subsequent work conducted by Saifullah et al.
+A subsequent work by Saifullah et al.
 (2020) made improvements by leveraging both local and global contexts.
 Huang et al.
-(2023a) leveraged pre-trained CodeBert (Feng et al., 2020), a transformer-based masked language model to predict FQNs in the code snippet.
-Compared to LLMs evaluated in this study, CodeBert is much smaller at 125 million parameters compared to even Llama3.1:8b’s 8 billion parameters.
+(2023a) employed the pre-trained CodeBert (Feng et al., 2020), a transformer-based masked language model to predict FQNs in the code snippet.
+Compared to LLMs evaluated in this study, CodeBert is much smaller with 125 million parameters, whereas even Llama3.1:8b’s contains 8 billion parameters.
+Chen et al.
+(2024) proposed a hybrid approach that iteratively combines constraint-based techniques, such as SnR, with machine learning methods to refine the results further.
 Kabir et al.
-(2024) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt.
+(2025) introduced ZS4C, an LLM-based approach that employs multiple prompts and incorporated compiler error messages to infer and correct compilation errors in the code snippet using LLMs.
 However, since our work focuses on mitigating data leakage, which fundamentally undermines inference reliability regardless of the prompting strategy used, we used a single prompt.
+More broadly, data leakage can also diminish the contribution of ZS4C’s use of compiler error messages, as it enables LLMs to achieve artificially high scores on StatType-SO, thereby masking the substantial improvements gained when compiler error messages are incorporated.
 8.2. Type Inference for Dynamically Typed Languages
@@
 Pradel et al.
-(2020) propose TypeWriter, which uses a deep learning-based approach to predict the types and then utilizes a search-based approach to validate the predicted type.
+(2020) propose TypeWriter, which uses a deep learning-based approach to predict the types and then utilizes a search-based approach to validate the predicted types.
 Peng et al.
-(2022) propose HiTyper, which is a deep learning-based approach combining static type inference.
+(2022) propose HiTyper, a deep learning-based approach that combines with static type inference.
 It conducts static inference and DL-based prediction iteratively to construct a type dependency graph, which records type dependency information among variables.
@@
 8.3. LLM Data Leakage
-Data leakage is not limited to type inference but also affects other software engineering tasks, such as automated program repair using LLMs (Xia et al., 2023; Ouyang et al., 2024).
+Data leakage is not limited to type inference but also affects other software engineering tasks.
+Currently, LLMs have been integrated into a wide variety of software engineering tools such as automated code repair (Yuan et al., 2022; Zhang et al., 2024; Xia et al., 2023; Xia and Zhang, 2024; Ouyang et al., 2024; Bouzenia et al., 2025; Ye et al., 2025; Kong et al., 2025b), code generation (Jiang et al., 2024a; Mu et al., 2024; Li et al., 2025; Fan et al., 2025), code refactoring (Shirafuji et al., 2023; Pomian et al., 2024), and code completion (Guo et al., 2023; Wu et al., 2024; Semenkin et al., 2025), which can all be potentially affected by data leakage either currently or in the future.
 Xia et al.
@@
 Sainz et al.
-examined several academic datasets and reported that most of them were either used as training data for ChatGPT or likely exposed to it (Sainz et al., 2023).
-A subsequent work done by Golchin and Surdeanu (2023) proposes a more advanced detection technique and detects the presence of test data of several datasets in the training data of LLMs.
+(2023b, a) examined several academic datasets and reported that most of them were either used as training data for ChatGPT or likely exposed to it.
+A subsequent work by Golchin and Surdeanu (2023) proposes a more advanced detection technique and detects the presence of test data of several datasets in the training data of LLMs.
+Marone and Van Durme (2023) propose data portraits to efficiently check whether a given string is present in the training data, assuming the training data is available.
+Recently, Kong et al.
+(2025a), detected memorization in LLMs using low-probability events such as exact code repair matches.
+However, for type inference on code snippets, there is generally one ground truth of types that are expected.
+Thus, exact matches do not precisely indicate data leakage.
 As LLM training datasets grow larger and increasingly opaque, detecting instances of data leakage during evaluation becomes progressively more challenging.
-The data leakage problem remains an open challenge for evaluating LLMs.
-Some studies have been conducted in other fields (Magar and Schwartz, 2022; Balloccu et al., 2024; Mirzadeh et al., 2024; Cao et al., 2024) regarding data leakage and LLM evaluation.
+Nevertheless, data leakage remains an open challenge for evaluating LLMs.
+Some studies have been conducted in other fields (Magar and Schwartz, 2022; Balloccu et al., 2024; Mirzadeh et al., 2024; Cao et al., 2024; Uddin et al., 2025; Elangovan et al., 2021) regarding data leakage and LLM evaluation.
 Mirzadeh et al.
-(2024) showed that LLMs have significant limitations when conducting genuine mathematical reasoning which suggests that LLMs are conducting sophisticated pattern matching rather than true logical reasoning.
-A recent study (Cao et al., 2024) demonstrated that the performance of LLMs is far behind undergraduate students on the proposed project-level Java benchmark that exercises OOP features.
+(2024) showed that LLMs have significant limitations when conducting genuine mathematical reasoning, which suggests that LLMs are conducting sophisticated pattern matching rather than true logical reasoning.
+A recent study by (Cao et al., 2024) demonstrated that the performance of LLMs is far behind undergraduate students on the proposed project-level Java benchmark that exercises object-oriented programming features.
 Our evaluation on type inference rather than math questions or code generation also showed a significant drop in performance with generated code snippets which suggests that understanding the type system in code is challenging for LLMs.
-So far, there is no conclusive evidence showing that LLMs are conducting true type inference on code snippets.
-Small changes to the inputs can drastically alter model outputs (Shi et al., 2023; Mirzadeh et al., 2024; Jiang et al., 2024).
+So far, there is no conclusive evidence showing that LLMs are conducting genuine type inference on code snippets.
+Small changes to the inputs can drastically alter model outputs (Shi et al., 2023; Mirzadeh et al., 2024; Jiang et al., 2024b).
 While some guidelines have been raised for software engineering research using LLMs (Ozkaya, 2023; Sallou et al., 2024), protecting against data leakage remains an open challenge.
@@
 This paper conducted a comprehensive assessment of LLMs’ type inference capabilities on Java code snippets.
-Our three-pronged evaluation mitigated potential data leakage issues and identified possible limitations of LLMs.
-First, we create a new dataset named ThaliaType.
-By evaluating the performance of LLMs on ThaliaType and StatType-SO, we conclude that LLM type inference performance may not generalize well to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall.
-In addition, LLM type inference performance diminishes when inferring less commonly used FQNs, potentially introducing bias and presenting opportunities for future research to enhance performance.
-Second, we designed and applied three semantic-preserving code transformations to code snippets in ThaliaType and StatType-SO to investigate LLMs’ semantic understanding.
-Through evaluating LLMs with these transformed code snippets, we find that the precision and recall of LLMs in type inference depend on the exact content of the code snippet, even when the semantic meaning of the code remains unchanged.
-This suggests that LLMs may rely heavily on superficial patterns in the code rather than truly understanding its semantics, harming LLMs’ performances’ generaliability.
-Third, we apply delta debugging to analyze the minimal syntactic elements sufficient for LLMs to infer types.
-We find that LLMs can still correctly infer FQNs even when the types were absent from the snippets, that StatType-SO may allow LLMs to rely on superficial patterns rather than deep code understanding.
-Our experiments suggest that LLMs’ strong performance in prior evaluations was likely due to data leakage rather than a genuine understanding and analysis of the semantics in code snippets.
-Our findings suggest that future evaluations of LLMs should utilize a generated dataset, carefully designed to assess the true capabilities of LLMs for type inference tasks.
-All of our code and dataset are available at https://github.com/uw-pluverse/thalia-type.
+First, to address potential data leakage we introduced a new benchmark suite, ThaliaType, and compared it against the widely used StatType-SO.
+Second, using StarCoder2 as a baseline (which we confirmed to suffer from data leakage), our evaluation revealed that all tested LLMs exhibited similar performance degradations on unseen code snippets.
+Specifically, we observed up to a 59% decrease in precision and a 72% decrease in recall.
+Moreover, LLM performance diminished when inferring less commonly used FQNs, suggesting potential bias and presenting opportunities for future research.
+Finally, we designed and applied three semantic-preserving code transformations to both benchmark suites, to investigate LLMs’ understanding of the execution semantics.
+While LLMs maintained consistent performance under simple transformations, their performance significantly declined when transformations are combined on StatType-SO.
+These results on StatType-SO, not observed on ThaliaType, indicate that LLM performance on StatType-SO may be inflated by data leakage and may not be generalizable.
+Overall, our findings underscore the need for future evaluations to incorporate unseen benchmarks, such as newly generated ThaliaType code snippets and transformed variants, while discontinuing StatType-SO.
+This approach provides a more reliable assessment of LLMs’ performance in type inference tasks.
+All code and datasets used in this paper are publicly available at https://github.com/uw-pluverse/thalia-type.